### v2

# Orak: A Foundational Benchmark for Training and Evaluating LLM Agents on Diverse Video Games

Abstract
Large Language Model (LLM) agents are reshaping the game industry, by enabling more intelligent and human-preferable characters. Yet, current game benchmarks fall short of practical needs: they lack evaluations of diverse LLM capabilities across various game genres, studies of agentic modules crucial for complex gameplay, and fine-tuning datasets to adapt pre-trained LLMs into gaming agents. To fill these gaps, we present Orak, a benchmark for training and evaluating LLM agents across 12 popular video games spanning all major genres. Using a plug-and-play interface built on Model Context Protocol (MCP), Orak supports systematic and reproducible studies of agentic modules in varied game scenarios. We further release a fine-tuning dataset of expert LLM gameplay trajectories covering multiple genres, turning general LLMs into effective game agents. Orak offers a united evaluation framework, including game leaderboards, LLM battle arenas, and in-depth analyses of input modality, agentic strategies, and fine-tuning effects, establishing a foundation towards versatile gaming agents. Code is available at https://github.com/krafton-ai/Orak.
1 Introduction
Large Language Model (LLM) agents are revolutionizing various industries (Wang et al., 2024a), and well-established benchmarks play a key role in unlocking their abilities on complex tasks, e.g., Coding (Hendrycks et al., 2021, Zhuo et al., 2024), Web Search (Liu et al., 2023, Pan et al., 2024, Levy et al., 2024), and Scientific Research (Mühlbacher et al., 2024, Zhang et al., 2025). In game industry, there is growing interest in using LLM agents to enhance user game experiences e.g., more intelligent non-player characters (NPCs), monsters, or companions (NVIDIA, 2025). In response, many benchmarks have been proposed to assess the ability of LLMs to play games (Hu et al., 2024b).
While these benchmarks have effectively utilized games to evaluate LLMs’ general capabilities, they exhibit three major limitations: 1) they often rely on text-only games or 2D-grid simulators rather than complex real video games, 2) they offer insufficient assessment of agentic modules, such as self-reflection, memory, and tool use, which are essential to complex gameplay, and 3) they lack fine-tuning datasets necessary to adapt pre-trained LLMs into effective gameplay agents, which significantly hinder the adoption of LLM agents in real-world video games.
To this end, we present Orak, a foundational benchmark designed to evaluate LLM agents across diverse video games. As shown in Figure 1, Orak includes 12 video games played by millions to billions of users worldwide: Street Fighter III, Super Mario, Ace Attorney, Her Story, Pokémon Red, Darkest Dungeon, Minecraft, Stardew Valley, StarCraft II, Slay the Spire, Baba Is You, and 2048. These games span 6 major game genres, i.e., action, adventure, role-playing, simulation, strategy, and puzzle, enabling a comprehensive assessment of key abilities required for versatile gameplay; action games enable testing fine-grained player control, adventure games challenge long-term memory and error handling, and strategy/puzzle games require complex logical reasoning and multi-step planning. With the use of real video games, Orak also ensures evaluation on rich, dynamic environments with varying stages, levels, and story-driven quests, which are even challenging for humans.
To enable consistent evaluation of rapidly evolving LLMs, we introduce a plug-and-play interface using Model Context Protocol (MCP) (Hou et al., 2025), allowing LLMs to seamlessly interact with agentic modules in gameplay. Each game environment and agentic module package operates as an independent MCP server, providing game mechanics (e.g., retrieving game states, executing game steps) or agentic strategies (e.g., reflection, planning) as callable tools to LLMs. During evaluation, LLM interacts with these servers by sequentially retrieving game states, performing action inference using agentic modules, and executing game steps, which enables streamlined evaluation across diverse games and supports controlled studies of various agentic modules.
In addition, we release a fine-tuning dataset that aims to transform pre-trained LLMs into effective gaming agents. The dataset consists of game interaction trajectories generated by expert LLMs, e.g., GPT-4o, using core agentic strategies on all games in Orak. These trajectories encapsulate meta-knowledge on how and when to use the agentic strategies to play various genres of games, leading to more resource-efficient and effective gaming agents.
Our benchmark offers comprehensive evaluation dimensions, including game score leaderboards, competitive LLM battle arenas, and in-depth analyses of visual input state, agentic strategies, and fine-tuning effects. Extensive experiments on Orak with 15 LLMs reveal that (1) proprietary LLMs achieve superior performance across games with significant gaps from open-source LLMs, (2) their performance gap becomes narrow in battle scenarios, (3) proprietary LLMs benefit from extended agentic workflows, while open-source LLMs show limited gains, (4) visual states often hinder gameplay performance yet, and (5) fine-tuning enables effective transfer of gameplay meta-knowledge from larger LLMs to smaller ones, leading to generalization in intra-game, out-of-distribution (OOD) game, and non-game unseen scenarios like math and web interaction. We believe that Orak not only establishes a foundation for building gaming agents but also serves as a critical benchmark for evaluating general LLMs on realistic, long-horizon decision-making tasks.
2 Related Work
Playing Games by LLMs.
Many works have explored the use of LLMs for gameplay. Early efforts focused on text-based games such as Jericho (Hausknecht et al., 2020), Zork (Tsai et al., 2023), and TextCraft (Prasad et al., 2023), where LLMs navigate textual environments with reasoning. Subsequent work shifted toward 2D-grid games, including Chess (Feng et al., 2023), NetHack (Küttler et al., 2020), and Crafter (Hafner, 2021), where spatial reasoning and puzzle-solving skills became more important. More recently, several studies have applied LLMs, combined with agentic workflows, to play more complex video games, such as Minecraft (Fan et al., 2022, Wang et al., 2023), Civilization (Qi et al., 2024), Pokémon (Hu et al., 2024c), and StarCraft (Ma et al., 2024). However, these approaches rely on manually customized agentic workflows for each specific game, limiting their usability toward developing a general gaming agent.
Evaluation Benchmarks for LLMs with Games.
As gameplay requires complex cognitive abilities, e.g., context understanding, logical reasoning, and error handling, several recent benchmarks have sought to evaluate LLMs or Vision Language Models (VLMs) on games (Hu et al., 2024b). GAMA-Bench (Huang et al., 2024), GameBench (Costarelli et al., 2024), GameArena (Hu et al., 2024a), and SmartPlay (Wu et al., 2023) focus on text-based games, assessing LLMs’ ability to navigate and reason on textual environments. Barlog (Paglieri et al., 2024) and LVLM-Playground (Wang et al., 2025) are mainly based on 2D-grid games, such as TicTacToe and Chess, to evaluate the spatial and visual reasoning capabilities of LLMs/VLMs. Using video games, Cradle (Tan et al., 2024) evaluates VLMs in 1 adventure and 3 simulation games, V-MAGE (Zheng et al., 2025) assesses VLMs on 5 action games, DSGBench (Tang et al., 2025) validates LLMs on 6 strategic games, and LMGame-Bench (Hu et al., 2025) mainly focus on puzzle games. Despite their contributions, existing benchmarks lack full coverage of game genres, omit in-depth studies on agentic modules, often rely on visual inputs, and under-explore the way to align pre-trained LLMs into versatile game agents. Table 1 summarizes the key characteristics of game benchmarks.
| Benchmarks | Game Domain | Full Genre | Games | Model Type | Agent Ablation | Fine-tuning Set |
| GAMA-bench (Huang et al., 2024) | Text | ✗ | 8 | LLM | ✗ | ✗ |
| GameBench (Costarelli et al., 2024) | Text | ✗ | 9 | LLM | ✗ | ✗ |
| GameArena (Hu et al., 2024a) | Text | ✗ | 6 | LLM | ✗ | ✗ |
| SmartPlay (Wu et al., 2023) | Text/2D-grid | ✗ | 6 | LLM | ✗ | ✗ |
| Balrog (Paglieri et al., 2024) | Text/2D-grid | ✗ | 6 | LLM/VLM | ✗ | ✗ |
| LVLM-Playground (Wang et al., 2025) | 2D-grid | ✗ | 6 | VLM | ✗ | ✗ |
| Cradle (Tan et al., 2024) | Video | ✗ | 4 | VLM | ✗ | ✗ |
| V-MAGE (Zheng et al., 2025) | Video | ✗ | 5 | VLM | ✗ | ✗ |
| DSGBench (Tang et al., 2025) | Video | ✗ | 6 | LLM | ✗ | ✗ |
| LMGame-Bench (Hu et al., 2025) | Video | ✗ | 6 | LLM/VLM | ✗ | ✗ |
| Orak (Ours) | Video | ✓ | 12 | LLM/VLM | ✓ | ✓ |
Fine-tuning LLMs toward Agents.
To enhance agent capability, many efforts have proposed agentic strategies, e.g., chain-of-thought reasoning (Wei et al., 2022, Yao et al., 2023), self-reflection (Shinn et al., 2023, Park et al., 2023), hierarchical task planning (Huang et al., 2022b; a), and skill libraries (Wang et al., 2023). Complementing these advancements, efforts have focused on fine-tuning strategies tailored to LLM agents, which span two main directions: data-centric approaches, which involve fine-tuning on curated expert demonstrations (Chen et al., 2023, Zeng et al., 2023, Chen et al., 2024); and framework-oriented approaches, which focus on learning from agentic interactions (Feng et al., 2024, Chen et al., 2025, Putta et al., 2024). Notably, FireAct (Chen et al., 2023) emphasizes unified data formatting and CodeAct (Wang et al., 2024b) highlights the importance of high-quality curation of training trajectories. However, fine-tuning methods for game-playing agents remain largely explored. Unlike structured tasks in web, programming, or math domains, games involve large, dynamic, and partially observable state spaces, which requires agents to generalize across a variety of situations and learn diverse behavior patterns, posing unique challenges.
3 Orak
We propose Orak, a benchmark designed to evaluate LLM agents across diverse video games. Figure 2 illustrates the evaluation pipeline of Orak, which is self-explanatory with code structures. By integrating the MCP interface with game environments and agentic modules, Orak enables systematic and plug-and-play evaluation of backbone LLMs with agentic strategies across various games. For evaluation, the game score is obtained by simply configuring the game, LLM backbone, and agentic strategy in eval.py. At each game step, the game observation is retrieved, the specified agent strategy is executed by the LLM, and the resulting action is applied to the game. This loop continues until the game ends or reaches the maximum step limit, after which the game score is recorded. Note that, with MCP interface, users can readily customize their agentic strategy, i.e., calling a single agentic module or multiple agentic modules sequentially. In the following section, we describe the 12 game environments in Orak, along with the submission guidelines for the benchmark leaderboard.
3.1 Game Environments
LLM Capabilities Required.
LLM agents require diverse capabilities to play games. Figure 3 summarizes the capability levels needed for each game in Orak, measured using principled criteria adopted from the context of game design (Wu et al., 2023, Koster, 2013):
Rule Following (RF). The level is measured by the extent to which adherence to rules is required for gameplay (1: single rule, 2: fewer than 5 rules, 3: 5 or more rules).
Logical Reasoning (LR). The number of LLM’s reasoning hops required to determine an in-game action (1: 1 hop, 2: 1 to 3 hops, 3: 3 or more hops).
Spatial Reasoning (SR). The level of spatial understanding required for gameplay (1: not necessary, 2: required in specific situations, 3: critical to core gameplay).
Long-text Understanding (LTU). The extent of long-context comprehension required for gameplay (1: a few lines, 2: a few paragraphs, 3: longer than one page with 500 words).
Long-term Planning (LP). The extent to which strategic planning is required (1: not necessary, 2: planning for up to 3 sequential actions, 3: essential to plan more than 3 sequential actions).
Error Handling (EH). The extent to which error correction is required during gameplay (1: not necessary, 2: requires a one-step rollback, 3: requires multi-step rollback and re-planning).
Odds Handling (OH). The extent to which understanding randomness is required for gameplay (1: not necessary, 2: randomness exists in game, 3: randomness is critical to core gameplay).
The capability level is measured on a scale of 1 to 3 by 8 human participants, and the moderate value is reported. Since most video games are designed to require various cognitive abilities even for humans, they tend to require high levels of LLM capabilities in many dimensions. For example, action games, Street Fighter III and Super Mario in red, require spatial reasoning and rule following, more than long-context understanding and planning, while adventure games, Ace Attorney and Her Story in yellow, emphasize long-text understanding and logical reasoning due to the need to comprehend long storylines. More detailed analysis of required LLM capabilities for each game is in Appendix B.
Game Description. For each game, we provide a brief description of (1) the game state, (2) the action space given to LLMs, and (3) the evaluation task and metric. More detailed explanations of each environment are elaborated in Appendices C-N.
(a) Street Fighter III (Capcom, 1997) is a 2D side-scrolling fighting action game with 20 unique characters, each with distinctive skills. (1) Game state: The player character, opponent character, remaining time, each player’s health, super-bar, stun-bar gauges, and the distance between the two characters. (2) Action space: 15-20 discrete actions: ‘move closer’, ‘low punch’, ‘high kick’, etc. (3) Evaluation task: Beating the game bot; performance is measured by the number of stages cleared.
(b) Super Mario (Christian Kauten, 2018) is a side-scrolling game where the player controls Mario to avoid obstacles, defeat enemies, and reach the flag. (1) Game state: The positions (x,y) and sizes of obstacles and enemies extracted from the current game state. (2) Action space: Mario keeps moving to the right, and LLM decides the jump level, discretized in 6 bins. (3) Evaluation task: Reaching out to the final flag; performance is measured by the horizontal distance that Mario travels before dying.
(c) Ace Attorney (Capcom, 2001) is a courtroom adventure game where players act as defense attorneys, gathering evidence and witnesses to prove their client’s innocence. (1) Game state: Dialogue history, collected evidence, court records profiles, etc. (2) Action space: Player’s courtroom actions: advancing dialogue, accessing court records, pressing witnesses, and presenting evidence. (3) Evaluation task: Performance is measured by response correctness and total steps taken.
(d) Her Story (Barlow, 2015) is an interactive adventure game where players explore police interview clips to uncover a hidden truth. (1) Game state: History of queries and search results with metadata for the first 5 clips (visual description, date, viewing status, and transcript if played). (2) Action space: Searching for clips with keywords, or selecting a video to play. (3) Evaluation task: Uncover the truth; Performance is measured by the number of distinct video clips viewed to complete the game.
(e) Pokémon Red (Game Freak, 1996) is a role-playing game where a player explores, collects Pokémon, and battles other trainers to progress the storyline. (1) Game state: Player’s location, party Pokémon (species, level, HP, status), inventory, battle state, and screen text. (2) Action space: Choosing high-level tools or low-level joypad actions. (3) Evaluation task: Defeat Brock, the first gym leader; Progress measured by how many of 12 predefined storyline flags are triggered.
(f) Darkest Dungeon (Red Hook Studios, 2016) is a role-playing game where heroes explore dungeons while managing stress and resources. (1) Game state: Party status (character stats, health, stress, and status effects), available skills, and enemy encounters. (2) Action space: Combat actions like ‘attack’, ‘heal’, and ‘swap’. (3) Evaluation task: Complete the first expedition; Performance is measured by the sum of the successful combats, the survived heroes, and their remaining stress.
(g) Minecraft (Mojang Studios, 2011) is an open-ended sandbox game where players explore a world, gather resources, and survive by placing and breaking blocks. (1) Game state: The player’s position, inventory, health status, the nearby blocks and biome, etc. (2) Action space: Executable JavaScript code within the Mineflayer environment (contributors, 2013). (3) Evaluation task: Crafting a target item; performance is measured by whether the item is collected in the inventory.
(h) Stardew Valley (ConcernedApe, 2016) is an open-ended life simulation game where players farm, fish, mine, and explore. (1) Game state: Player’s location, energy, inventory, crop status, soil status, date, and weather. (2) Action space: Leaving the house, entering the house, sleeping, buying seeds, tilling soil, watering, harvesting, and selling crops. (3) Evaluation task: Earning the most money by harvesting crops within the first 13 in-game days; performance is measured by total profit.
(i) StarCraft II (Blizzard Entertainment, 2010) is a real-time strategy game where players gather resources, construct buildings, train units, and command armies to defeat opponents. (1) Game state: Resource levels, unit/building counts, production queues, research progress, and observed enemy info. (2) Action space: 72 discrete actions, including unit training, building, research, and strategic operations. (3) Evaluation task: Beating built-in AI bots; performance is measured by the win rate.
(j) Slay the Spire (MegaCrit, 2017) is a deck-building roguelike game where players ascend a multi-floor tower, battling enemies and building decks. (1) Game state: Player’s class, deck, hand, health, relics, energy, enemies’ intents and statuses, and current floor. (2) Action space: Playing a card during combat, ending the turn, and selecting a card reward after combat. (3) Evaluation task: Defeating the final boss at the top floor; performance is measured by the number of floors reached.
(k) Baba Is You (Hempuli, 2019) is a puzzle game where a player manipulates the rules by moving word tiles on a board. (1) Game state: Coordinates of text and object tiles, and active rules. (2) Action space: A single movement ‘up’, ‘down’, ‘left’, and ‘right’, or a sequence of such moves. (3) Evaluation task: Solving the first stage; if the stage is not cleared, partial credit is awarded based on sub-goals (e.g., breaking the ‘Wall Is Stop’ rule).
(l) 2048 (Cirulli, 2014) is a sliding tile puzzle game that aims to combine numbered tiles on a 44 grid board to create a tile of the value 2048. (1) Game state: The current configuration of the 44 grid, where each cell contains either a number (power of 2) or is empty. (2) Action space: Four discrete actions; ‘up’, ‘down’, ‘left’, and ‘right’. (3) Evaluation task: Creating the 2048 tile; performance is measured by the normalized progress toward creating the 2048 tile.
3.2 Submission Guideline to Benchmark Leaderboard
Models. Participants can submit new pre-trained LLMs, VLMs, or their fine-tuned versions. Unless otherwise specified, all models will be evaluated under the default agentic strategies for each game specified in Section 5.1.
Agentic Strategies. Participants can also submit new agentic modules and strategies. All submitted strategies should follow a consistent structure across games, e.g., by maintaining a fixed sequence such as reasoning, planning, and using a specific tool.
4 Fine-tuning: Aligning Pre-trained LLMs into Game Agents
| Data example from Reflection module | Data example from Action module | ||||||||||||
We collect the fine-tuning dataset from expert LLMs, e.g., GPT-4o and o3-mini, playing all 12 games in Orak using several agentic modules. This dataset with environment interaction trajectories encapsulates meta-knowledge on how to use the agentic strategies to solve diverse game genres.
Data Format. We denote LLMs’ gameplay trajectory as , where is the number of game steps, and denotes the sequence of LLM inferences executed by agentic strategies at game step . Each LLM inference sequence is represented as , where is the -th agentic module in sequence, is the prompt for agentic module , is the game state, and is the corresponding response of LLM. Table 2 shows detailed data examples of .
Data Selection. For each game in Orak, we collect gameplay trajectories of expert LLMs, until we have more than 1000 LLM inference sequence in all collected. To ensure high-quality data, we sort the collected in terms of the game score and select the trajectories with the highest game scores until the number of selected exceeds 300. All selected trajectories follow the ‘reflection-planning-action’ sequence, so that we have around 300 samples for each agent module. By performing data selection on all 12 games, our fine-tuning set consists of approximately 11k samples.
Data Augmentation. To enhance the linguistic diversity, we augment each data sample by paraphrasing. We prompt GPT-4o to rephrase the game prompt while preserving all game-related information, generating 10 augmented samples for each sample . See Appendix O for prompting details and the effect of the number of augmentations.
Our fine-tuning dataset is mainly for supervised fine-tuning (SFT). While dynamic data extraction from the environment could enable reinforcement learning fine-tuning, we leave this for future work.
5 Experiment
5.1 Experiment Setup
Models.
We validate the performance of 15 LLMs using states provided in text format. The models include 8 open-source LLMs: LLaMA-3.2-1B/3B (Grattafiori et al., 2024), LLaMA-3.3-72B (Grattafiori et al., 2024), Qwen-2.5-3B/7B/72B (Yang et al., 2024), and Minitron-4B/8B (Sreenivas et al., 2024), and 7 proprietary LLMs: GPT-5/4o/4o-mini (Achiam et al., 2023), o3-mini, Gemini-2.5-pro (Team et al., 2023), Claude-3.7-sonnet (Anthropic, 2025), and DeepSeek-R1 (Guo et al., 2025). In addition, we study the effects of incorporating image inputs on 5 multi-modal LLMs: Qwen2.5-vl-7B/32B (Bai et al., 2025), GPT-4o, Gemini-2.5-pro, and Claude-3.7-sonnet.
Default Agentic Strategies.
For each game in Orak, we select the most effective agentic strategy over GPT-4o as the default agent strategy. Specifically, For Street Fighter III, HerStory, Darkest Dungeon, and 2048, we use a ‘zero-shot’ action inference agent. For Super Mario, Pokémon Red, Stardew Valley, StarCraft II, Slay the Spire, and Baba Is You, we use a ‘reflection-planning’ agent that sequentially performs self-reflection, subtask planning, and action inference, integrated with memory at each game step. For AceAttorney, we use a ‘reflection’ agent. For Minecraft, we use a ‘skill-management’ agent that further includes knowledge retrieval and skill management in the reflection-planning-action agent, following (Wang et al., 2023).
Metrics and Implementation Details.
For each game, we report the normalization score rather than the absolute score, i.e., the game score is normalized by the maximum game score. We report the average score of 3 to 20 trials for each game. We also report the average score of 3 human novices. More detailed metrics and LLM hyperparameter configurations can be found in Appendices C-N.
5.2 LLM Gameplay Performance
| Genre | Action | Adventure | RPG | Simulation | Strategy | Puzzle | Avg | ||||||
| Games | SF3 | SuperMario | AceAttorney | HerStory | Pokémon | DarkestD | Minecraft | Stardew | StarCraft2 | SlaySpire | BabaIsYou | 2048 | Rank |
| Llama-3.2-1B | 0.00.0 | 18.78.6 | 1.32.2 | 2.11.2 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 6.711.5 | 0.00.1 | 13.5 |
| Llama-3.2-3B | 13.35.8 | 31.810.1 | 4.61.3 | 4.21.1 | 0.00.0 | 47.539.2 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.30.2 | 11.4 |
| Qwen-2.5-3B | 20.00.0 | 23.414.1 | 20.017.4 | 1.21.1 | 0.00.0 | 44.822.2 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 13.311.5 | 0.10.1 | 12.1 |
| Qwen-2.5-7B | 16.711.5 | 27.29.6 | 9.30.2 | 8.52.0 | 0.00.0 | 88.82.0 | 0.00.0 | 0.00.0 | 0.00.0 | 5.00.0 | 20.00.0 | 0.60.4 | 10.5 |
| Minitron-4B | 16.711.5 | 24.46.0 | 35.74.5 | 4.62.3 | 0.00.0 | 0.00 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.10.0 | 11.6 |
| Minitron-8B | 23.35.8 | 31.312.8 | 29.93.6 | 8.21.8 | 0.00.0 | 63.830.4 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.70.7 | 10.1 |
| Llama3.3-70B | 33.328.9 | 32.011.8 | 53.923.4 | 46.410.8 | 16.714.4 | 87.28.2 | 0.00.0 | 40.937.2 | 0.00.0 | 5.00.0 | 20.00.0 | 1.41.3 | 7.3 |
| Qwen2.5-72B | 26.75.8 | 29.98.7 | 14.98.5 | 41.01.7 | 27.89.6 | 84.06.6 | 0.00.0 | 18.016.7 | 0.00.0 | 9.05.3 | 20.00.0 | 0.20.1 | 8.8 |
| GPT-4o-mini | 16.711.5 | 28.88.8 | 28.42.8 | 21.25.6 | 0.00.0 | 81.35.8 | 46.07.0 | 16.127.8 | 75.050.0 | 3.32.9 | 13.311.5 | 1.11.0 | 9.2 |
| GPT-4o | 29.714.3 | 34.114.2 | 85.31.5 | 64.25.2 | 38.99.6 | 93.41.5 | 71.07.0 | 81.44.8 | 100.00.0 | 23.622.1 | 20.00.0 | 5.61.5 | 3.6 |
| GPT-5 | 19.511.0 | 31.610.5 | 59.126.5 | 74.91.7 | 88.94.8 | 92.90.7 | 73.07.0 | 92.38.6 | 25.050.0 | 26.219.4 | 100.00.0 | 10.22.1 | 3.6 |
| o3-mini | 33.315.3 | 34.914.6 | 91.71.5 | 66.53.6 | 0.00.0 | 89.02.1 | 75.00.0 | 55.116.0 | 25.050.0 | 15.00.0 | 73.346.2 | 25.37.3 | 4.0 |
| Gemini-2.5-pro | 13.311.5 | 38.014.4 | 55.73.4 | 67.53.3 | 83.30.0 | 93.71.6 | 75.00.0 | 59.210.1 | 100.00.0 | 51.931.9 | 73.346.2 | 5.12.5 | 3.5 |
| Claude-3.7 | 16.711.5 | 31.78.2 | 81.91.6 | 62.92.6 | 63.919.2 | 89.92.5 | 75.00.0 | 53.620.9 | 50.057.7 | 15.00.0 | 46.746.2 | 5.32.7 | 5.1 |
| Deepseek-R1 | 20.00.0 | 28.713.2 | 83.31.5 | 67.23.9 | 75.00.0 | 91.71.1 | 41.70.0 | 66.111.5 | 50.057.7 | 24.917.1 | 20.00.0 | 11.53.4 | 5.1 |
| Human (Novice) | 20.020.0 | 100.00.0 | 87.82.6 | 72.62.8 | 86.112.7 | 90.32.0 | 70.80.0 | 69.819.9 | 33.357.7 | 52.923.2 | 100.00.0 | 22.710.4 | - |
Table 3 shows the gameplay performance of LLMs on Orak with default agentic strategies. Overall, proprietary LLMs outperform open-source LLMs across all games in most cases. Gemini-2.5-pro performs the best on average, ranking first in 5 out of 12 games, with the best average ranking of 3.5. gpt-5/4o forms the second-best group with an average ranking of 3.6. GPT-5 shows superiority in puzzle games, i.e., Baba Is You and 2048, which require strong mathematical and logical reasoning, and spatial understanding abilities. Most small open-source LLMs, size under 8B, show almost zero score on complex games, i.e., Pokémon-red, Minecraft, Stardew Valley, StarCraft II, and Slay the Spire. Mid-sized open-source models, Llama3.3-70B and Qwen2.5-72B, show advances in moderately hard games, such as Stardew Valley and Slay the Spire, but their performance remains far below that of proprietary LLMs. See Appendices C-N for more detailed results on each game.
5.3 LLM Arena
Among 12 games in Orak, Street Fighter III and StarCraft II support two-player competitive modes. For Street Fighter III, we conduct pairwise battles among 8 LLMs with ‘zero-shot’ agent. Each pair competed in three rounds, and the agent winning 2 out of 3 rounds was declared the winner. To ensure a fair comparison, both agents were assigned the same character, Ken, in all matches. Figure 4(a) shows the relative win rates and Elo ratings. Interestingly, different from the result in Section 5.2, Minitron-8B consistently outperforms all other LLMs and shows the best Elo rating. This may imply that when multiple agents are involved in the environment, adversarial actions can change the game dynamics. For StarCraft II, we conduct pairwise battles among 7 LLMs, with each pair competing in a single round. Both agents were assigned the same race, Protoss, in all matches. As in Figure 4(b), Claude-3.7-Sonnet shows the best Elo rating, while GPT-4o and Minitron-8B form the second group.
5.4 Ablation Study for Agentic Modules
| Models | Agent | Action | Adventure | RPG | Simulation | Strategy | Puzzle | Avg | ||||||
| Strategies | SF3 | SuperMario | AceAttorney | HerStory | Pokémon | DarkestD | Minecraft | Stardew | StarCraft2 | SlaySpire | BabaIsYou | 2048 | Rank | |
| LLaMA-3B | Zeroshot | 13.35.8 | 21.28.2 | 5.73.2 | 4.21.1 | 0.00.0 | 47.539.2 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.30.2 | 6.4 |
| Reflection | 30.017.3 | 32.48.6 | 4.61.3 | 4.41.3 | 0.00.0 | 47.339.0 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.00.0 | 5.6 | |
| Planning | 20.00.0 | 27.08.4 | 4.61.3 | 5.21.0 | 0.00.0 | 56.323.6 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.10.1 | 6.2 | |
| Ref-Plan | 16.720.8 | 31.810.1 | 3.80.0 | 5.40.4 | 0.00.0 | 57.031.6 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.10.2 | 6.1 | |
| GPT-4o | Zeroshot | 29.714.3 | 29.69.2 | 49.91.3 | 64.25.2 | 33.30.0 | 93.41.5 | 0.00.0 | 34.529.9 | 50.057.7 | 24.79.2 | 20.00.0 | 5.61.5 | 3.3 |
| Reflection | 23.320.8 | 32.315.5 | 85.31.5 | 61.50.4 | 36.14.8 | 85.210.3 | 50.00.0 | 15.64.5 | 0.00.0 | 41.319.6 | 20.00.0 | 3.52.9 | 3.0 | |
| Planning | 30.026.5 | 29.412.5 | 52.70.5 | 59.45.7 | 33.30.0 | 82.08.6 | 13.00.0 | 54.920.2 | 50.057.7 | 35.39.2 | 20.00.0 | 6.05.5 | 3.1 | |
| Ref-Plan | 23.320.8 | 34.114.2 | 52.80.5 | 62.04.9 | 38.99.6 | 91.62.5 | 50.00.0 | 81.44.8 | 100.00.0 | 36.025.5 | 20.00.0 | 7.05.7 | 2.2 | |
Table 4 shows the ablation results of LLaMA-3.2-3B and GPT-4o across 4 agent strategies. Interestingly, the impact of adding agentic modules to gameplay performance differs between the two LLMs. For GPT-4o, the inclusion of agentic modules consistently improves gameplay performance; ‘reflection-planning’ agent achieves the best average ranking of 2.2, followed by ‘reflection’ and ‘planning’ agents with rankings of 3.1 and 3.3, and ‘zero-shot’ agent with the lowest ranking of 3.4. However, LLaMA-3.2-3B does not follow this trend. The ‘reflection’ agent shows the highest average ranking of 5.6, while ‘reflection-planning’ follows with 6.1, although it adds the planning module. This indicates that, for relatively smaller LLMs like LLaMA-3.2-3B, adding agentic modules may increase the complexity of the prompt, hindering their decision-making accuracy. These results suggest that the optimal agentic strategy may depend on the inherent capability of the LLM.
5.5 Effect of Visual Input
Setup.
To study the effect of visual input, we divide games into two groups: Group 1 (Table 6) includes games where the provided textual game state can entirely be derived from a visual screenshot. Group 2 (Table 6) comprises games where the textual state contains information beyond what is visible in the current frame (e.g., abstracted inventory lists or off-screen character/item details); for this group, we exclude evaluation on Image-only input types for fairness. Minecraft was excluded since the environment does not provide the visionary view of a bot that LLM plays.
Result.
As shown in Table 6, relying solely on Image-only input led to a substantial drop in performance. This was consistently reflected across all models, with average ranks deteriorating significantly compared to Text-only input. In contrast, as shown in Table 6& 6, utilizing Both text and visual inputs produced mixed effects on game scores and model ranks. For instance, in games like Street Fighter III (Group 1), since on-screen visual details are challenging to fully convey textually, adding visual context significantly benefited Claude, increasing its score by . Conversely, in narrative-heavy games such as Ace Attorney (Group 2), the same approach often proved detrimental; GPT-4o’s score, for example, dropped by , and its average rank in that group fell from 2.9 (Text-only) to 5.0. This highlights that the impact of combining modalities varies considerably, with some scenarios showing improved ranks or scores while others demonstrated a decline.
| Models | Input | SF3 | SuperMario | Stardew | BabaIsYou | 2048 | Rank |
| Qwen2.5-7B | Text | 0.00.0 | 26.08.2 | 0.00.0 | 13.311.5 | 0.10.1 | 12.8 |
| Image | 3.35.8 | 25.110.1 | 0.00.0 | 0.00.0 | 0.20.3 | 13.5 | |
| Both | 3.35.8 | 25.68.0 | 0.00.0 | 0.00.0 | 0.40.5 | 12.7 | |
| Qwen2.5-32B | Text | 16.75.8 | 32.010.2 | 15.222.7 | 20.00.0 | 0.20.3 | 4.3 |
| Image | 33.320.8 | 25.810.7 | 0.0 0.0 | 13.311.5 | 0.40.3 | 5.1 | |
| Both | 33.315.3 | 29.76.6 | 26.913.9 | 20.00.0 | 0.20.3 | 3.4 | |
| GPT-4o | Text | 29.714.3 | 34.114.1 | 81.44.8 | 20.00.0 | 5.61.5 | 1.7 |
| Image | 23.715.9 | 27.113.7 | 0.0 0.0 | 6.711.5 | 1.81.1 | 5.3 | |
| Both | 24.314.5 | 27.110.2 | 41.922.1 | 20.00.0 | 5.44.5 | 3.0 | |
| Gemini-2.5 | Text | 13.311.5 | 38.013.4 | 59.210.1 | 73.346.2 | 5.12.5 | 5.2 |
| Image | 16.711.5 | 28.510.7 | 7.69.0 | 20.00.0 | 5.52.4 | 7.4 | |
| Both | 20.010.0 | 40.99.6 | 60.06.0 | 86.723.1 | 3.12.6 | 4.0 | |
| Claude-3.7 | Text | 16.711.5 | 28.713.2 | 53.620.9 | 46.746.2 | 5.32.7 | 5.8 |
| Image | 23.311.5 | 25.66.4 | 0.00.0 | 20.00.0 | 8.44.0 | 8.0 | |
| Both | 33.35.8 | 22.66.3 | 49.81.0 | 20.00.0 | 6.70.9 | 6.2 |
| Models | Input | AceAttorney | HerStory | Pokémon | DarkestD | StarCraft2 | SlaySpire | Rank |
| Qwen2.5-7B | Text | 10.00.0 | 3.21.5 | 2.84.8 | 82.71.5 | 0.00.0 | 0.00.0 | 9.1 |
| Both | 17.613.1 | 2.50.8 | 2.84.8 | 81.83.0 | 0.00.0 | 0.00.0 | 9.3 | |
| Qwen2.5-32B | Text | 71.921.5 | 15.42.8 | 27.89.6 | 89.91.5 | 0.00.0 | 0.00.0 | 7.1 |
| Both | 68.610.0 | 17.04.6 | 25.014.4 | 91.03.0 | 0.00.0 | 0.00.0 | 7.0 | |
| GPT-4o | Text | 85.31.5 | 64.25.2 | 38.99.6 | 93.41.5 | 100.00.0 | 23.622.1 | 2.9 |
| Both | 53.51.7 | 40.629.5 | 41.78.3 | 92.23.0 | 50.057.7 | 23.622.1 | 5.0 | |
| Gemini-2.5 | Text | 55.73.4 | 67.53.3 | 83.30.0 | 93.71.6 | 100.00.0 | 51.931.9 | 2.1 |
| Both | 52.60.8 | 64.92.4 | 83.30.0 | 92.21.8 | 100.00.0 | 26.219.4 | 3.2 | |
| Claude-3.7 | Text | 81.91.6 | 62.92.6 | 63.919.2 | 89.92.5 | 50.057.7 | 15.00.0 | 4.8 |
| Both | 71.317.3 | 63.63.1 | 72.24.7 | 90.15.7 | 50.057.7 | 9.74.6 | 4.3 |
5.6 Effect of Fine-tuning
| Models | Finetune | Intra-Game | OOD-Game | Non-Game | |||||||
| SF3 | DarkestD | StarCraft2 | SlaySpire | BabaIsYou | SuperMario | 2048 | Math500 | WebShop-E | WebShop-H | ||
| Llama-3.2-1B | ✗ | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 18.78.6 | 0.10.1 | 24.11.6 | 2.70.5 | 1.90.8 |
| ✓ | 42.016.4 | 93.42.6 | 0.00.0 | 8.03.5 | 20.00.0 | 26.712.3 | 2.81.8 | 25.61.7 | 0.50.0 | 0.00.0 | |
| Llama-3.2-3B | ✗ | 12.011.0 | 87.29.5 | 0.00.0 | 0.00.0 | 20.00.0 | 31.810.1 | 0.10.2 | 29.61.8 | 0.00.0 | 0.00.0 |
| ✓ | 40.07.1 | 92.00.2 | 0.00.0 | 10.71.2 | 20.00.0 | 34.47.0 | 3.12.5 | 36.10.9 | 8.40.3 | 12.60.8 | |
Setup.
To study the effect of fine-tuning, we consider three types of generalization: Intra-game, OOD-game, and Non-game. Intra-game generalization evaluates whether an LLM can adapt to unseen scenarios within the same game, e.g., new stages or characters. OOD-game generalization evaluates whether fine-tuning on a specific set of games enables the model to act better on unseen games. Non-game generalization evaluates whether our fine-tuning set with agent gameplay trajectory helps the model to perform better on non-game tasks like math or web navigation. For Intra-game and Non-game generalizations, we fine-tune LLMs on all 12 games, while for OOD-game generalization, we split 12 games in to 10 training games and 2 test games, i.e., Super Mario and 2048. See Appendix P for detailed fine-tuning configurations and unseen scenario details in intra-game generalization.
Intra-game Generalization.
As shown in Table 7, fine-tuned Llama-3.2-1B/3B show better generalization to unseen scenarios than the pretrained ones; in 3 out of 5 games, both fine-tuned models outperform their pretrained counterparts. The performance gain largely comes from the model learning to generate valid actions more reliably after fine-tuning, especially in environments where the pretrained model frequently failed to act meaningfully. However, this approach appears insufficient to significantly enhance the spatial reasoning abilities of smaller models. This is evident in Baba Is You, where fine-tuned models struggle to construct the winning condition required to score above 20.0.
OOD-game Generalization.
Similarly, despite being trained exclusively on trajectories from different games, fine-tuned models perform significantly better on OOD games like Super Mario and 2048. This suggests that fine-tuning on trajectories shaped by reflection and planning enables models to learn transferable decision-making routines, as the underlying capabilities required for these behaviors are shared across games. Specifically, the improvement in 2048 likely benefits from its structural similarity to Baba Is You, a training game that also uses 2D grid layout and discrete move actions (up, down, left, right). More fine-tuning analysis with varying OOD gaps are in Appendix Q.
Non-game Generalization.
More interestingly, LLMs fine-tuned on our gameplay trajectories sometimes show improved performance in non-game tasks like Math500 and WebShop (Yao et al., 2022). In Table 7, LLaMA-3.2-3B consistently improves on both Math500 and Webshop. In particular, on WebShop-E (easy) and -H (hard), it shows a dramatic gain from 0.0 to 8.4 or 12.6, indicating our fine-tuning set is more effective for agentic tasks that require decision-making abilities similar to gameplay. In contrast, the smaller model, Llama-3.2-1B, shows no observable generalization to non-game tasks, which suggests that the extent of generalization may depend on model capacity.
6 Discussion
Conclusion.
In this paper, we introduce Orak, a benchmark designed to train and evaluate LLM agents across diverse video games. Orak enables comprehensive assessments of LLM capabilities required to play most game genres. Through a plug-and-play interface powered by MCP, it allows consistent evaluation of rapidly evolving LLMs over various agentic modules. In addition, we release a fine-tuning dataset of game interaction trajectories of top-performing LLMs, which can effectively transform pre-trained LLMs into gaming agents. With the comprehensive game set and user-friendly interface, Orak sets a new foundation for game-based LLM evaluation, driving progress towards versatile and high-performing gaming agents.
Limitations.
We note that Orak provides environments favorable for LLM reasoning and decision-making by pre-processing game states into structured text where information unnecessary to gameplay is hidden. This may offer insights for providers seeking to deploy resource-efficient LLMs on games. We leave a complementary direction, providing the full in-game states with rich uncurated texts to LLMs, for future work. Further discussions, including cost and licensing, are provided in Appendix R.
Appendix A Game Genre Categorization
Video games are generally categorized into six widely recognized genres, i.e., Action, Adventure, Role-Playing, Simulation, Strategy, and Puzzle, each characterized by distinct gameplay structures (Wikipedia contributors, 2025). Action games emphasize responsiveness and precise physical control. Adventure games focus on narrative exploration, interactive dialogues, and clue-based progression grounded in logical inference. Role-Playing games focus on character progression, stat-driven combat mechanics, and quest-based narrative development. Simulation games present system-driven environments in which players manage complex and interdependent variables such as time, resources, and procedural systems. Strategy games are a genre that emphasizes planning, resource management, decision-making, and tactical execution. Puzzle games revolve around rule-based problem solving, pattern recognition, and logical or spatial reasoning, typically within a clearly defined system of constraints.
In Table 8, we categorize each of the 12 video games in Orak into one of the six major genres based on the primary gameplay characteristics and the genre information provided by the respective game publishers. Some games belong to one or more genres due to hybrid gameplay mechanics.
| Game | Action | Adventure | Role-Playing | Simulation | Strategy | Puzzle |
| Street Fighter III | ||||||
| Super Mario | ||||||
| Ace Attorney | ||||||
| Her Story | ||||||
| Pokémon Red | ||||||
| Darkest Dungeon | ||||||
| Minecraft | ||||||
| Stardew Valley | ||||||
| StarCraft II | ||||||
| Slay the Spire | ||||||
| Baba Is You | ||||||
| 2048 |
(a) Street Fighter III (Capcom, 1997) is classified as an Action game, as it primarily relies on responsiveness, frame-precise input, and physical dexterity. Its gameplay requires fast reflexes and mastery of complex input sequences.
(b) Super Mario (Christian Kauten, 2018) is categorized as Action game. The core gameplay emphasizes precise timing in jumping and movement, demanding moment-to-moment control in response to environmental hazards and enemy placements.
(c) Ace Attorney (Capcom, 2001) is labeled as an Adventure game due to its narrative-driven structure, reliance on clue collection, and logical deduction. Players progress by interacting with characters and uncovering story elements through investigative mechanics, with minimal emphasis on reflex-based input.
(d) Her Story (Barlow, 2015) similarly fits within the Adventure category. Though more experimental in form, it shares a strong focus on narrative discovery through a search-based interface, requiring players to piece together a fragmented story using non-linear exploration and deductive reasoning.
(e) Pokémon Red (Game Freak, 1996) is classified as a Role-Playing game, as its primary mechanics involve turn-based combat, character progression, stat management, and quest-driven exploration. Additionally, Adventure was assigned as a secondary genre to reflect the game’s emphasis on world exploration, interaction with non-player characters, and sequential progression through narrative landmarks.
(f) Darkest Dungeon (Red Hook Studios, 2016) is similarly assigned Role-Playing as the main genre, supported by stat-driven character development and progression systems. However, it also includes significant Strategy elements, as players must carefully manage resources, form party compositions, and make tactical decisions in turn-based combat. Adventure was further added as a secondary genre due to its dungeon-crawling structure and emphasis on risk-driven exploration.
(g) Minecraft (Mojang Studios, 2011) is categorized as a Simulation game due to its open-ended, system-driven mechanics, including resource gathering, crafting, and environmental manipulation. It also exhibits Role-Playing traits through its progression systems and player-driven narrative development, warranting secondary classification.
(h) Stardew Valley (ConcernedApe, 2016) is also assigned to the Simulation genre based on its emphasis on time management, farming systems, and interrelated mechanics spanning multiple in-game variables. It incorporates Role-Playing through relationship-building and character progression, and Adventure through dungeon exploration, seasonal events, and quest-based interactions.
(i) StarCraft II (Blizzard Entertainment, 2010) is classified as a Strategy game, consistent with its real-time strategic planning, resource allocation, and micromanagement mechanics. Given the importance of unit-level control, Action was added as a secondary genre to reflect the real-time, reflex-driven demands during gameplay.
(j) Slay the Spire (MegaCrit, 2017) is categorized solely as a Strategy game. The gameplay centers around deck-building, route optimization, and turn-based combat, requiring players to plan several moves.
(k) Baba Is You (Hempuli, 2019) is clearly identified as a Puzzle game, as its mechanics are centered on solving logic-based problems through the manipulation of in-game rules represented by words. The core loop involves constrained, rule-based problem solving and spatial reasoning.
(l) 2048 (Cirulli, 2014) is also classified as a Puzzle game, characterized by deterministic mechanics, arithmetic pattern recognition, and constraint-based spatial logic within a fixed grid.
This genre classification provides a structured foundation for analyzing agent cognition and gameplay dynamics across diverse games.
Appendix B Required LLM Capabilities for Gameplay
Looking at the Figure 3, most games demand advanced LLM capabilities across multiple dimensions, as they are designed to challenge a wide range of human cognitive skills. (1) Action games, Street Fighter III and Super Mario in red, require spatial reasoning and rule following, more than long-context understanding and planning. (2) Adventure games, Ace Attorney and Her Story in yellow, emphasize long-text understanding and logical reasoning due to the need to comprehend long storylines. (3) Role-playing games, Pokémon Red and Darkest Dungeon in brown, require strong long-term planning, logical reasoning, and rule-following abilities to understand game-specific rules and complete milestones of game tasks. (4) Simulation games, Minecraft and Stardew Valley in green, also require high levels of long-term planning and rule-following abilities. While Minecraft requires strong spatial reasoning and error handling, which are generally essential for simulation games, Stardew Valley gets lower scores for them because these abilities are not critical for its evaluation task; earning money by harvesting crops. (5) Strategic games, StarCraft II and Slay the Spire in blue, require various LLM abilities for gameplay. Notably, these two games are the only ones in Orak that require 5 different LLM capabilities rated at level 3, highlighting that recent strategic video games increasingly demand a wide range of cognitive skills for effective gameplay. (6) Puzzle games, Baba Is You and 2048 in purple, require high levels of spatial reasoning, logical reasoning, and long-term planning because puzzle games are typically designed to require complex problem-solving through multiple reasoning hops and spatial understanding.
Appendix C Street Fighter III
C.1 Game Description for Street Fighter III
(a) Ken. (b) Chun-Li.
Environment. Street Fighter III (Capcom, 1997) is a 2D competitive fighting game, known for precise controls, deep mechanics, and a diverse roster of characters. Each character features unique moves, combos, and super arts, requiring precise timing and strategic decision-making. Players aim to defeat their opponent through a mix of normal attacks, special moves, and advanced mechanics like parries and cancels. For implementation, we use Diambra Arena environment (DIAMBRA, 2025), a Docker-based platform designed for RL research. Street Fighter III is one of the environments supported by Diambra, which not only enables seamless extraction of the game state—such as screenshots, health, timer, and super bar values—but also provides a straightforward interface for sending controller inputs. In this setting, the agent operates in a discrete action space that directly corresponds to various controller actions, including directional movement, attack buttons, and their combinations, enabling intuitive and fine-grained control of the in-game character.
The game supports both single-player and multi-player modes. In single-player mode, the player progresses through ten increasingly difficult stages, facing stronger opponents at each level. Each stage follows a best-of-three format, where the player must win two out of three matches to advance to the next stage. The game ends upon either completion of the final stage or defeat. In multi-player mode, two players compete in a best-of-three match, and the game is over once a winner is determined. For our default evaluation setting, agents play Ken in both modes (see Figure 5(a)). However, since the environment supports a variety of characters, we also conduct evaluations using Chun-Li to demonstrate intra-game generalization capabilities (see Figure 5(b)).
Observation-to-Text Conversion. The Diambra environment offers a convenient interface for extracting the game state from Street Fighter III. Through this interface, we obtain the latest game frame at a resolution of 224×384, along with key state information such as remaining time, player and opponent health, super bar gauge, super count, stun bar gauge, and stun status. However, a critical aspect in fighting games is understanding the relative positions of the characters, which is not directly provided by the Diambra environment. To address this limitation, we employ a lightweight YOLOv11 object detection model (Khanam & Hussain, 2024) to extract the relative positions of the two characters from the game frame. Based on the computed distance, we classify the spatial relationship into three discrete categories—very close, close, and far—and incorporate this information into the agent’s user prompt.
Action Space. In the Diambra environment, the native action space is defined on a per-frame basis and consists of 18 discrete actions. These are composed of:
Idle action (1 total): Idle (No action)
Movement actions (8 total): Left, LeftUp, Up, UpRight, Right, RightDown, Down, DownLeft
Attack actions (9 total): Low Punch, Medium Punch, High Punch, Low Kick, Medium Kick, High Kick, Low PunchLow Kick, Medium PunchMedium Kick, High PunchHigh Kick
To enable more strategic and temporally consistent behavior, we use a higher-level action space that abstracts these frame-level controls into semantically meaningful commands. Each high-level action is mapped to a predefined sequence of low-level controller inputs, often spanning multiple frames. The high-level action space is divided into two categories:
Character-agnostic actions (14 total): Move Closer, Move Away, Jump Closer, Jump Away, Super Attack, Low Punch, Medium Punch, High Punch, Low Kick, Medium Kick, High Kick, Low PunchLow Kick, Medium PunchMedium Kick, High PunchHigh Kick
Character-specific actions:
Ken: Fireball (Hadouken), Hurricane Kick, etc. -
Chun-Li: Kikkoken, Hyakuretsukyaku, etc.
For example, if the character is positioned on the left side of the screen and the high-level action is ‘Move Closer’, the system issues ‘Right’ movement commands over four frames. If the action is ‘Fireball’, the corresponding low-level sequence would be ‘Down’‘DownRight’‘Right’‘Medium Punch’. Since the number of character-specific actions varies, the total size of the high-level action space differs depending on character, typically ranging around 20 actions.
C.2 Gameplay Prompt for Street Fighter III
Our implementation of Street Fighter III supports four types of agents: reflection, planning, reflection-planning, and zero-shot. Among these, we introduce prompts for the ‘reflection-planning’ agent in this subsection. Figures 7–9 present the prompts used by the reflection-planning agent for planning, action inference, and reflection, respectively. At each step, the agent plans to determine the subtask, using the planning module. Based on this subtask, the action inference module infers the optimal action to execute. Finally, the reflection module evaluates whether the executed action was successful.
Planning prompt. As shown in Figure 7, the system prompt provides detailed instructions for an agent to support strategic planning. It defines the assistant’s role in proposing suitable subtasks based on the target task and the current game state. The user prompt includes: (1) the main goal of the game, (2) the previous subtask (generated by the recent planning module), (3) the last executed action, (4) a self-reflection on the last action (generated by the recent reflection module), (5) the current state, and (6) the expected output format for the subtask reasoning task.
Action inference prompt.
As shown in Figure 8, the system prompt outlines strategic guidelines for playing Ken, a predefined set of valid actions, and the required output format.
The user prompt contains: (1) the current subtask (provided by the recent planning module), (2) the last executed action, (3) the corresponding self-reflection (generated by the recent reflection module), (4) the current state, and (5) the expected output format for the action inference task.
Reflection prompt. As illustrated in Figure 9, the system prompt provides detailed instructions for an agent to perform reflection. The agent is required to analyze whether the last action was successful based on state transitions. The user prompt includes: (1) the target task, (2) the current subtask (generated by the recent planning module), (3) the last executed action, (4) the previous state, (5) the current state, and (6) the expected output format for the reflection task.
C.3 Evaluation Metric for Street Fighter III
Single-Agent Play. In the single-player mode, the agent faces a series of 10 stages against in-game rule-based bots. The game ends either when the player loses a stage or successfully clears all 10 stages. Therefore, the evaluation metric can be straightforwardly defined as
Multi-Agent Play. To evaluate models in a competitive multi-agent environment, we conduct pairwise matches between all agents and compute Elo ratings based on their win rates. For each pair, three games are played to obtain a reliable estimate of relative performance. The resulting win rate matrix and Elo scores are presented in Figure 4(a).
We adopt a Bradley-Terry model formulation (Bradley & Terry, 1952) for Elo estimation, where each model’s rating is iteratively optimized using gradient ascent on the log-likelihood of observed outcomes. The gradient is computed based on the expected win probabilities derived from current ratings, using the standard Elo transformation:
where and denote the Elo ratings of agent and agent , respectively. The expected probability of defeating increases as the rating difference becomes larger. After optimization, we shift all Elo ratings so that their mean equals 1,500 for intuitive interpretation. If two models receive identical ratings, the one that won their head-to-head match is ranked higher.
C.4 Experimental Configuration for Street Fighter III
For all 6 open-source LLMs, including Llama-3.2-1B/3B, Qwen-2.5-3B/7B, and Minitron-4B/8B, we use a temperature of 0.0 and a repetition penalty of 1.0. During LLM inference, we pause the Street Fighter III game environment. We set the maximum number of game steps to 10,000. Due to the in-game time constraints, episodes do not reach the maximum of 10,000 steps. A single stage (best-of-three matches) is usually resolved within 100 to 200 steps, as rounds either timeout or end early when one player’s health reaches zero.
C.5 Result for Street Fighter III
All reported results in Tables 11–12 are computed as the mean and standard deviation over three to five independent runs, using a ‘zero-shot’ agent for each model configuration. To establish a baseline for comparison, we additionally included a random agent that selects an arbitrary action uniformly at random. This agent was evaluated over 30 episodes, and its performance is summarized in tables.
Single-Agent Play. The performance across various LLMs was evaluated as presented in Table 11. The smallest model, Llama-3.2-1B, completely failed to comprehend the current game context and consistently ignore the required output format, performing even worse than a random agent. On the other hand, all other evaluated models surpassed the random agent, demonstrating varying levels of competence. Commercial LLMs generally outperformed their open-source counterparts, with GPT-4o and o3-mini standing out as notable examples.
To investigate the effectiveness of agentic modules, we conducted an ablation study shown in Table 11. For Llama-3.2-3B, the Reflection agent showed notably superior performance, indicating that reflective reasoning significantly aids in aligning actions to game dynamics, possibly by allowing the model to reassess and correct previous outputs based on feedback. Conversely, for GPT-4o, both the zero-shot and Planning agents performed remarkably well. This may suggest GPT-4o’s inherent capability for generalization (zero-shot) and structured sequential reasoning (planning), enabling efficient decision-making without iterative reflection.
Input modality may significantly influence agent performance, since spatial information in fighting games is expecially important for gameplay. Since our implementation simplify character distance into three sparse levels, we expected image inputs would enhance spatial understanding and consequently improve performance. However, as shown in Table 11, GPT-4o surprisingly demonstrated decreased performance when using image inputs, possibly due to limited visual comprehension capabilities. In contrast, Gemini and Claude effectively leveraged visual data, improving their gameplay scores.
Lastly, Table 12 demonstrates intra-game generalization capabilities when fine-tuning models on Ken-specific gameplay data and evaluating on Chun-Li scenarios, which represent out-of-distribution conditions due to differences in action spaces (as previously detailed in Section C.1). Remarkably, the previously format-incompliant pretrained Llama-3.2-1B learned to follow the required output format effectively and exhibited excellent gameplay performance after fine-tuning, even in Chun-Li gameplay. A similar improvement was observed in Llama-3.2-3B, which even surpassed GPT-4o in performance. These results highlight the significant impact of targeted fine-tuning, demonstrating that relatively small-scale LLMs can achieve substantial gains in task-specific capability.
| Models | SF3 | Rank |
| Random Agent | 10.06.4 | 12 |
| Llama-3.2-1B | 0.00.0 | 13 |
| Llama-3.2-3B | 13.35.8 | 10.5 |
| Qwen-2.5-3B | 20.00.0 | 4.5 |
| Qwen-2.5-7B | 16.711.5 | 7.5 |
| Minitron-4B | 16.711.5 | 7.5 |
| Minitron-8B | 23.35.8 | 3 |
| GPT-4o-mini | 16.711.5 | 7.5 |
| GPT-4o | 29.714.3 | 2 |
| o3-mini | 33.315.3 | 1 |
| Gemini-2.5-pro | 13.311.5 | 10.5 |
| Claude-3.7 | 16.711.5 | 7.5 |
| Deepseek-R1 | 20.00.0 | 4.5 |
| Models | Agent | SF3 | Rank |
| Random Agent | - | 10.06.4 | 9 |
| Llama-3B | Zero-shot | 13.35.8 | 8 |
| Reflection | 30.017.3 | 1.5 | |
| Planning | 20.00.0 | 6 | |
| Ref-Plan | 16.720.8 | 7 | |
| GPT-4o | Zero-shot | 29.714.3 | 3 |
| Reflection | 23.320.8 | 4.5 | |
| Planning | 30.026.5 | 1.5 | |
| Ref-Plan | 23.320.8 | 4.5 |
| Models | Input | SF3 | Rank |
| Random Agent | - | 10.06.4 | 10 |
| GPT-4o | Text | 29.714.3 | 2 |
| Image | 23.715.9 | 4 | |
| Both | 24.314.5 | 3 | |
| Gemini | Text | 13.311.5 | 9 |
| Image | 16.711.5 | 7 | |
| Both | 20.010.0 | 6 | |
| Claude | Text | 16.711.5 | 7 |
| Image | 23.311.5 | 5 | |
| Both | 33.35.8 | 1 |
| Model | Finetune | SF3 |
| Llama-3.2-1B | ✗ | 0.00.0 |
| ✓ | 42.016.4 | |
| Llama-3.2-3B | ✗ | 12.011.0 |
| ✓ | 40.07.1 | |
| GPT-4o | ✗ | 10.00.0 |
Multi-Agent Play. We evaluate agent performances in a multi-agent environment by conducting pairwise matches among 8 LLMs, all operating in a zero-shot setting. The results are provided in Figure 4(a). Each pairwise matchup consisted of three independent games, with each game played in a best-of-three format to determine the winner. To ensure fair performance comparison, both agents used the same character—Ken—in all matches.
Notably, unlike the single-agent evaluation results in Table 11, Minitron-8B consistently outperforms all other models in the multi-agent arena and achieves the highest Elo rating. This divergence raises the possibility that the involvement of other intelligent agents could alter the game dynamics, perhaps due to increased strategic diversity or emergent adversarial behavior.
Appendix D Super Mario
D.1 Game Description for Super Mario
(a) Screenshot. (b) Object Patterns.
Environment. Super Mario (1985 Super Mario Bros) (Christian Kauten, 2018) is a side-scrolling game where the player controls Mario to avoid obstacles, defeat monsters, and reach the flag. In this environment, Mario progresses through the game using directional key controls (e.g., ‘left’ and ‘right’ keys) and jump actions. Mario should either destroy or traverse obstacles (e.g., bricks, stairs, pipes), avoid or defeat monsters (e.g., Goombas, Koopas) by jumping on them and avoid falling into pits. For implementation, we adopt the gym_super_mario_bros environment (Kauten, 2018), which is widely used in the reinforcement learning (RL) community. Specifically, we use the SuperMarioBros-v1 environment, where Mario plays within a 256240-pixel screen with a black background, as shown in Figure 10(a). The environment consists of 8 worlds, each containing 4 stages. We use World 1, Stage 1 as our default evaluation setting (for Table 3). However, the environment supports evaluation across any world and stage. For example, we use World 3, Stage 1 for the evaluation of intra-game generalization (for Table 7).
Observation-to-Text Conversion. The environment only provides RGB image frames as observations. To convert visual game states into text input suitable for LLMs, we apply visual pattern matching to parse the exact location of each object on the frame. As shown in Figure 10(b), in the SuperMarioBros-v1 environment, the pixel-level visual patterns of objects remain stable across frames. By maintaining a set of pixel templates for all game objects as game assets, we perform 2D visual pattern matching to parse the presence and exact location of each object in the scene. These parsed object locations are converted to 2D coordinates and formatted as text, which is then passed to the LLM as part of its observation input.
Action Space. We constrain Mario to only move in the ‘right’ direction to simplify the control space. Mario can ‘jump’ at varying heights, and by constraining action spans 4 game frames (i.e., frame skipping), Mario’s jumping ability is discretized into 7 levels. Jump Level 0 corresponds to walking forward without jumping, while Jump Levels 1 through 6 represent increasing jump heights, with Level 6 being the highest possible jump. At each game step, the LLM chooses a jump level from 0 to 6, determining the jump height as Mario moves to the right.
D.2 Gameplay Prompt for Super Mario
Figure 11 shows the action inference prompt used by the ‘zero-shot’ agent for playing Super Mario. The system prompt contains most of the gameplay-specific knowledge. It includes (1) the main goal of the game, (2) detailed descriptions and sizes of each object, (3) explanations and safety notes for each action, and (4) the expected input-output format between the LLM and the environment. The user prompt provides the current game state as a list of all objects detected in the frame, represented by their top-left corner (x, y) coordinates obtained by visual pattern matching. Given this prompt, the LLM agent infers the appropriate jump level to advance safely toward the flag by avoiding obstacles and monsters.
D.3 Evaluation Metric for Super Mario
The goal of Super Mario is to reach the flag located at the right end of the stage. Since the gym_super_mario_bros environment provides Mario’s current position on the map, we define the evaluation metric as the proportion of the distance traversed toward the flag before Mario dies. Formally, the normalized score is defined as:
where , , and are the coordinate of Mario traversed before die, that of the flag, and that of the starting position on the map, respectively.
D.4 Experimental Configuration for Super Mario
For all 6 open-source LLMs, including Llama-3.2-1B/3B, Qwen-2.5-3B/7B, and Minitron-4B/8B, we use a temperature of 1 and a repetition penalty of 1. During LLM inference, we pause the Super Mario game environment. We set the maximum number of game steps to 100. We run all experiments with 20 trials and report the average score with the standard deviation.
D.5 Result for Super Mario
| Models | SuperMario | Rank |
| Llama-3.2-1B | 18.78.6 | 12 |
| Llama-3.2-3B | 31.810.1 | 4 |
| Qwen-2.5-3B | 23.414.1 | 11 |
| Qwen-2.5-7B | 27.29.6 | 9 |
| Minitron-4B | 24.46.0 | 10 |
| Minitron-8B | 31.312.8 | 6 |
| GPT-4o-mini | 28.88.8 | 7 |
| GPT-4o | 34.114.2 | 3 |
| o3-mini | 34.914.6 | 2 |
| Gemini-2.5-pro | 38.014.6 | 1 |
| Claude-3.7 | 31.78.2 | 5 |
| Deepseek-R1 | 28.713.2 | 8 |
| Models | Agent | SuperMario | Rank |
| Llama-3B | Zero-shot | 21.28.2 | 8 |
| Reflection | 32.48.6 | 2 | |
| Planning | 27.08.4 | 7 | |
| Ref-Plan | 31.810.1 | 4 | |
| GPT-4o | Zero-shot | 29.69.2 | 5 |
| Reflection | 32.315.5 | 3 | |
| Planning | 29.412.5 | 6 | |
| Ref-Plan | 34.114.2 | 1 |
| Models | Input | SuperMario | Rank |
| GPT-4o | Text | 34.114.1 | 3 |
| Image | 27.113.7 | 6.5 | |
| Both | 27.110.2 | 6.5 | |
| Gemini | Text | 38.013.4 | 2 |
| Image | 28.510.7 | 5 | |
| Both | 40.99.6 | 1 | |
| Claude | Text | 28.713.2 | 4 |
| Image | 25.66.4 | 8 | |
| Both | 22.66.3 | 9 |
As shown in Table 15, Gemini-2.5-pro achieves the highest score of 38.0 on Super Mario, followed by o3-mini with the score of 34.9. Among open-source LLMs, Llama-3.2-3B and Minitron-8B perform competitively, achieving scores of 31.8 and 31.3 respectively, which are comparable to the 31.7 score of Claude-3.7-sonnet. Qualitatively, most LLMs, including Gemini-2.5-pro, frequently fail to correctly estimate the parabolic jump trajectory when Mario is mid-air. This change of moves often results in Mario colliding with obstacles or being killed by monsters.
Table 15 shows the effect of reflection and planning modules on gameplay performance. Among these, the reflection module has a more pronounced impact. Specifically, when Mario is stuck in front of high obstacles such as warp pipes, the reflection module enables the agent to revise its previous low jump level decisions and select a higher jump level, allowing it to overcome the obstacle and proceed, thereby improving the final score. GPT-4o achieves the best score of 34.1 when both reflection and planning modules are used. In contrast, Llama-3.2-3B performs best when only the reflection module is used with a score of 32.4, indicating that Llama-3.2-3B may not benefit from the additional planning module or be easily disturbed by its response.
As shown in Table 15, using image-only input observations consistently underperforms compared to text-only inputs across all models, including GPT-4o, Gemini-2.5-pro, and Claude-3.7-sonnet. This suggests that relying solely on visual input makes it more challenging for models to extract detailed information from the game scene or to perform spatial reasoning, i.e., estimating distances between objects. In contrast, when both text and image inputs are provided, GPT-4o and Claude-3.7-sonnet show a performance drop, while Gemini-2.5-pro shows improved performance. This indicates that multimodal input can be beneficial when the model effectively integrates complementary information from both modalities.
Appendix E Ace Attorney
E.1 Game Description for Ace Attorney
Environment. Ace Attorney (Capcom, 2001) is a courtroom adventure game where players act as defense attorneys, gather evidence, and cross-examine witnesses. We target the first episode of Phoenix Wright: Ace Attorney Trilogy on Steam (see Figure 12). We define four subtasks: one three-question multiple-choice quiz where the player selects the correct answer, and three cross-examination tasks where the player presses witnesses for more details or presents evidence to expose contradictions. We use Harmony (Pardeike, 2025) with a BepInEx plugin (Contributors, 2025a) to hook the game’s source code at launch, capturing states such as dialogue text, arrow-button visibility, keyboard inputs, and Court Record entries. The hooks save states as .txt or .json files and monitor a command .txt file for inputs, injecting them into the game in real time.
Observation-to-Text Conversion. All states remain in text form and require only minimal post‐processing. We map speaker indices to character names (e.g., index ‘2’ ‘Phoenix Wright’) using a predefined mapping and replace original names with arbitrary aliases to prevent contamination (e.g., ‘Phoenix Wright’ ‘Alias’). We also convert Court Records and multiple‐choice candidate options—originally stored in .json—into continuous descriptive text when they exist.
Action Space. The basic actions include pressing ‘Ok’ to progress the dialogue and ‘Tab’ to access the Court Records. For multiple-choice questions, the action space expands to include candidate options. During cross-examinations, additional actions become available: ‘Left’ to return to the previous dialogue, ‘Press’ to question the witness further, selecting relevant evidence from the Court Records, and ‘Present’ to introduce it during the examination. Each option carries an index, and the model returns only the corresponding integer.
E.2 Gameplay Prompt for Ace Attorney
Figure 13 presents the zero-shot agent’s action-inference prompt for Ace Attorney. The system prompt specifies (1) the game’s goal, (2) procedures for dialogue, multiple-choice questions, and cross-examinations, (3) rules for accessing Court Records and presenting evidence, and (4) the expected I/O format between the LLM and the environment. The user prompt lists recent conversations—highlighting the conversation currently visible on the screen—and provides timestamped Court Record entries. Using this prompt, the agent detects contradictions, selects the correct actions, and manages dialogue and evidence to advance the trial.
E.3 Evaluation Metric for Ace Attorney
Each subtask begins and ends at fixed points, with screenshots at both start and end and the preceding conversation history provided, and yields a reward and a step count . In the multiple‐choice task (MC), is the number of correct answers out of three, while in cross‐examination tasks CE1, CE2, and CE3, is 1 for a pass and 0 for a fail; any failure incurs a maximum step count of 50. We normalize each reward and step count against fixed benchmarks and , namely for MC, for CE1, for CE2, and for CE3, by computing
and assign each task a difficulty weight reflecting MC, CE1, CE2, and CE3 respectively. These weights were set proportional to each task’s difficulty and the minimum number of steps required (the benchmarks). We then form weighted averages
and combine them into a composite score
with (70% for accuracy, 30% for efficiency). All experiments were repeated three times, and we report the sample mean and standard deviation of score.
E.4 Experimental Configuration for Ace Attorney
For all six open-source LLMs (Llama-3.2-1B/3B, Qwen-2.5-3B/7B, and Minitron-4B/8B), we set the temperature to 0.7, apply a repetition penalty of 1, cap game steps at 50, and limit conversation history to the most recent 20 exchanges.
E.5 Result for Ace Attorney
| Models | AceAttorney | Rank |
| Llama-3.2-1B | 1.32.2 | 12 |
| Llama-3.2-3B | 4.61.3 | 11 |
| Qwen-2.5-3B | 20.017.4 | 9 |
| Qwen-2.5-7B | 9.30.2 | 10 |
| Minitron-4B | 35.74.5 | 6 |
| Minitron-8B | 29.93.6 | 7 |
| GPT-4o-mini | 28.42.8 | 8 |
| GPT-4o | 85.31.5 | 2 |
| o3-mini | 91.71.5 | 1 |
| Gemini-2.5-pro | 55.73.4 | 5 |
| Claude-3.7 | 81.91.6 | 4 |
| Deepseek-R1 | 83.31.5 | 3 |
| Models | Agent | AceAttorney | Rank |
| Llama-3B | Zero-shot | 5.73.2 | 5 |
| Reflection | 4.61.3 | 6.5 | |
| Planning | 4.61.3 | 6.5 | |
| Ref-Plan | 3.80.0 | 8 | |
| GPT-4o | Zero-shot | 49.91.3 | 4 |
| Reflection | 85.31.5 | 1 | |
| Planning | 52.70.5 | 3 | |
| Ref-Plan | 52.80.5 | 2 |
| Models | Input | AceAttorney | Rank |
| GPT-4o | Text | 85.31.5 | 1 |
| Both | 53.51.7 | 5 | |
| Gemini | Text | 55.73.4 | 4 |
| Both | 52.60.8 | 6 | |
| Claude | Text | 81.91.6 | 2 |
| Both | 71.317.3 | 3 |
As shown in Table 18, o3-mini tops the leaderboard with a score of 91.7, followed by GPT-4o with a score of 85.3, Deepseek-R1 with a score of 83.3, and Claude-3.7-sonnet with a score of 81.9. Among open-source models, Gemini-2.5-pro and Minitron-4B score 55.7 and 35.7, respectively, while smaller Llama and Qwen variants remain below 20. Notably, Qwen-2.5-3B completes the first cross-examination in one out of three trials, where other open-source models struggle. Models scoring around 30 often misinterpret key details and repeat the same mistakes until they reach the maximum number of steps. Models below 20 sometimes reason correctly but mostly select irrelevant options, whereas the lowest performers fail to follow the required format and cannot advance further.
Table 18 compares our reflection and planning modules. The base Llama-3.2-3B model scores 5.7; adding reflection or planning alone yields approximately 4.6, and combining both drops the score to 3.8. Smaller models like Llama-3.2-3B often focus on incorrect details and produce flawed reasoning when reflection or planning is faulty. GPT-4o scores highest with reflection alone (85.3), since reflection prevents repeated errors and helps spot contradictions. Planning alone adds little, and combining it with reflection lowers the score to approximately 52.7, as it attempts to resolve contradictions not visible on-screen.
Table 18 compares input modalities. GPT-4o and Gemini-2.5-pro score 85.3 vs. 53.5 and 55.7 vs. 52.6 for text-only vs. multimodal inputs, respectively, while Claude-3.7-sonnet falls from 81.9 to 71.3. Because we supply complete text descriptions of every on-screen element and its context, adding visual input brings no improvement—confirming that text alone suffices to play Ace Attorney.
Appendix F Her Story
F.1 Game Description for Her Story
Environment. Her Story (Barlow, 2015) is an interactive adventure game where players explore police interview clips to uncover a hidden truth. The player begins the game by accessing an old desktop interface, where a program called L.O.G.I.C. Database is open. The player can enter keywords into the database to retrieve up to five video clips whose transcripts contain the searched word. By watching these clips and gathering clues, the player repeatedly formulates new queries to reconstruct the underlying story. To interface the game with our code, we use Harmony (Pardeike, 2025) together with Unity Doorstop (Contributors, 2025b) to log internal game states to a .txt file. Specifically, each line in the file is a JSON object representing a snapshot of the game’s state at a given moment. Each object contains an event type and its associated metadata. The following examples illustrate key elements of the logged game state. For clarity, we omit auxiliary metadata that are present in the actual logs but are only used for debugging or UI-related purposes, such as video IDs, screen resolution, and UI element positions.
Load title screen: {”status”: ”title”}
Load L.O.G.I.C. Database: {”status”: ”start_game”}
Query: {”status”: ”query”, ”keyword”: keyword}
Get query result: {”status”: ”query_result”, ”num_total”: number of clips containing the keyword, ”num_visible”: number of clips shown, ”video”: list of {”new”: 1 if not viewed, otherwise 0, ”session”: recording date, ”outfit”: visual description of the thumbnail}}
Open video panel: {”status”: ”open_detail”}
Close video panel: {”status”: ”close_detail”}
Play video: {”status”: ”play_video”, ”script”: transcript of the video}
Close video: {”status”: ”close_video”}
Observation-to-Text Conversion. We aggregate the game states and convert them into textual observations. Each observation includes summary information such as the number of clips containing the keyword and the number of clips shown. It also contains per-clip metadata, including the recording date, thumbnail description, viewing status, and the transcript if the clip was viewed after the query.
Action Space. The original Her Story game is designed as a point-and-click interface, where the player interacts with the game by typing keywords into a search bar, clicking on retrieved clip thumbnails, navigating panels, and controlling playback. These interactions rely on low-level input mechanisms such as mouse movements and keyboard input. To reduce complexity, we abstract these low-level interactions into two high-level actions: searching with a keyword and playing the retrieved video clip.
Search [keyword]: Returns all video clips whose transcripts contain the exact word. It consists of three low-level GUI actions: (1) clicking the search bar, (2) typing the keyword, and (3) pressing Enter to submit the query.
Play Video [i]: Plays the -th video from the current search result list. It consists of four low-level GUI actions: (1) clicking the thumbnail of the video clip to open the panel, (2) clicking the thumbnail within the panel to start playback, (3) either waiting until the video finishes or pressing the ESC key to exit playback, and (4) clicking the Exit button to close the panel.
F.2 Gameplay Prompt for Her Story
Figure 14 shows the action inference prompts used by the ‘zero-shot’ agent to play Her Story. The system prompt provides instructions covering: (1) the main goal of the game, (2) the type of information each video clip may contain, (3) behavioral rules the agent should follow—such as avoiding repeated keywords in the search history—and (4) the expected input-output format between the LLM and the environment. The user prompt includes: (1) the last executed action, (2) the current game state, and (3) the search history.
F.3 Evaluation Metric for Her Story
To uncover the truth behind the case, the player actively explores the video archive by issuing queries and watching clips. We evaluate this behavior by counting how many unique clips the player has viewed from the full archive of 271 clips. This aligns with the original game design, where unlocking achievements and reaching the ending depend on the number of clips watched. Specifically, we define the score as:
F.4 Experimental Configuration for Her Story
For all LLMs, we use a temperature of 0.3 and a repetition penalty of 1. We limit the maximum number of interactions with the game environment to 400 steps. We run each experiment three times and report the average score along with the standard deviation.
F.5 Result for Her Story
Table 21 summarizes the gameplay scores of different LLMs on Her Story. Commercial LLMs outperform open-source models in this task. Gemini-2.5-pro achieves the highest score of 67.3, followed by Deepseek-R1 (66.9), o3-mini (66.3), GPT-4o (64.0), Claude-3.7-sonnet (62.6), and GPT-4o-mini (21.1). The low score of GPT-4o-mini is due to its repeated search of keywords such as alibi, evidence, and witness, even after those queries have already been issued. This redundancy reduces its ability to discover new clips. Among open-source LLMs, all models score below 10. Most exhibited similar failure patterns: they fail to play unseen clips, repeatedly issue the same keywords, or search using entire sentences (e.g., for the name ”Simon”, for keywords related to Luna’s) rather than meaningful single words.
Table 21 reports the effect of incorporating reflection and planning modules on gameplay performance. For Llama-3.2-3B, the ‘reflection-planning’ agent achieves the best performance, followed by ‘planning’, ‘reflection’, and ‘zero-shot’ agents. This suggests that such modules can improve performance for weaker base models. In contrast, for GPT-4o, the ‘zero-shot’ agent achieves the highest score, and adding reflection or planning modules does not lead to further improvements. These modules tend to produce information that may be redundant for stronger models like GPT-4o, which can already generate effective queries without additional reasoning support.
Table 21 compares gameplay performance across different input modalities. In our multimodal setup, the visual input corresponds to the main screen of the L.O.G.I.C. Database interface, as shown in Figure 14. For GPT-4o and Gemini-2.5-pro, using both text and image inputs underperforms compared to the text-only setting. In particular, for GPT-4o, performance varies significantly depending on the random seed when using multimodal inputs. We observe issues such as repeated use of the same keywords and the inclusion of quotation marks around keywords (e.g., "murder"), which often lead to failed queries. In contrast, Claude-3.7-sonnet shows slightly better performance in the multimodal setting. We speculate two reasons why the visual input fails to improve performance in most cases: (1) most of the useful visual elements (e.g., clip lists, thumbnails) are already represented in the text observation, and (2) additional information that visual input could provide—such as the interviewee’s gestures or expressions—is only available when the video is actually played, and thus not present in the static visual input used in our setup.
| Models | HerStory | Rank |
| Llama-3.2-1B | 2.11.2 | 11 |
| Llama-3.2-3B | 4.21.1 | 10 |
| Qwen-2.5-3B | 1.21.1 | 12 |
| Qwen-2.5-7B | 8.52.0 | 7 |
| Minitron-4B | 4.62.3 | 9 |
| Minitron-8B | 8.21.8 | 8 |
| GPT-4o-mini | 21.25.6 | 6 |
| GPT-4o | 64.25.2 | 4 |
| o3-mini | 66.53.6 | 3 |
| Gemini-2.5-pro | 67.53.3 | 1 |
| Claude-3.7 | 62.92.6 | 5 |
| Deepseek-R1 | 67.23.9 | 2 |
| Models | Agent | HerStory | Rank |
| Llama-3B | Zero-shot | 4.21.1 | 8 |
| Reflection | 4.41.3 | 7 | |
| Planning | 5.21.0 | 6 | |
| Ref-Plan | 5.40.4 | 5 | |
| GPT-4o | Zero-shot | 64.25.2 | 1 |
| Reflection | 61.50.4 | 3 | |
| Planning | 59.45.7 | 4 | |
| Ref-Plan | 62.04.9 | 2 |
| Models | Input | HerStory | Rank |
| GPT-4o | Text | 64.25.2 | 3 |
| Both | 40.629.5 | 6 | |
| Gemini | Text | 67.53.3 | 1 |
| Both | 64.92.4 | 2 | |
| Claude | Text | 62.92.6 | 5 |
| Both | 63.63.1 | 4 |
Appendix G Pokémon Red
G.1 Game Description for Pokémon Red
Environment. Pokémon Red (Game Freak, 1996) is a role-playing game where the player navigates the Kanto region to catch and train creatures called Pokémon, battle other trainers, and ultimately defeat the Elite Four and the Champion. The player explores various environments, including towns, routes, caves, and buildings, encountering wild Pokémon and other characters. The gameplay loop involves exploring these areas, engaging in turn-based battles with Pokémon, and managing a team of up to six Pokémon. For implementation, we utilize the PyBoy (Baekalfen, ) emulator to run the game. Specifically, our evaluation focuses on a segment where the player starts in Pallet Town and progresses towards Viridian City, encountering wild Pokémon and trainers. The game screen resolution is 160144 pixels.
Observation-to-Text Conversion. Instead of relying on visual pattern matching, we directly access the game’s internal memory via the PyBoy emulator to extract relevant game state information. This includes detailed map information, the player’s current coordinates, information about the player’s party (e.g., Pokémon, their HP), encountered opponent Pokémon information, and the player’s inventory of items. This rich set of information is then formatted as text to serve as the observation input for the LLM.
Action Space. The fundamental action space consists of the Game Boy buttons: ‘up’, ‘down’, ‘left’, ‘right’, ‘a’, ‘b’, ‘start’, and ‘select’. Additionally, we define a set of higher-level tools to facilitate more complex interactions:
move_to(x, y): Finds and executes a path to the specified map coordinates .
interact_with_object(object_name): Interacts with a specified object in the environment.
warp_with_warp_point(warp_point_coord): Uses a specified warp point to move to a different location.
overworld_map_transition(direction): Transitions to an adjacent map in the given direction.
continue_dialog(): Advances the current dialogue.
select_move_in_battle(move_name): Selects and uses a specific move in a Pokémon battle.
switch_pkmn_in_battle(pkmn_name): Switches to a different Pokémon in the player’s party during a battle.
run_away(): Attempts to flee from a wild Pokémon battle.
use_item_in_battle(item_name): Uses a specified item during a Pokémon battle.
At each step, the LLM can choose up to five consecutive fundamental actions or invoke one of the provided tools.
G.2 Gameplay Prompt for Pokémon Red
Figure 15 shows the system prompt used by the agent for playing Pokémon Red, which provides the LLM with the necessary game rules, action space details (including basic controls and available tools), information about different game states, and the expected input/output format. The system prompt guides the LLM on how to interpret the game state and decide on the next action or tool to use to achieve the overarching goals of becoming the Champion and completing the Pokédex.
Figure 16 illustrates an example of the user prompt provided to the LLM. This includes the recent history of actions and their outcomes, the current game state (map information, player position, inventory, party, screen text, etc.), any recent critique on the agent’s actions, the current sub-task (if any), and relevant memory entries. Based on this information, the LLM infers the next action or tool to use, following the guidelines set in the system prompt.
G.3 Evaluation Metric for Pokémon Red
The goal in our defined segment of Pokémon Red is to progress through a series of key storyline milestones. We define 12 predefined storyline flags, and our evaluation metric is the percentage of these flags achieved by the agent. The 12 flags are: Exit Red’s House, Encounter Professor Oak, Choose a starter Pokémon, Finish the first battle with the Rival, Arrive in Viridian City, Receive Oak’s parcel, Deliver Oak’s parcel to Professor Oak, Obtain the Town Map, Purchase a Poké Ball, Catch a new Pokémon, Arrive in Pewter City, Defeat Pewter Gym Leader Brock. The final score is calculated as the percentage of these 12 flags that have been successfully triggered within a given episode or evaluation period. Formally,
G.4 Experimental Configuration for Pokémon Red
We configure all six open-source LLMs (Llama-3.2-1B/3B, Qwen-2.5-3B/7B, and Minitron-4B/8B) with a temperature of 0.1, a repetition penalty of 1, and a maximum of 1000 game steps. We conduct all experiments over three trials and report the mean score along with its standard deviation.
G.5 Result for Pokémon Red
Table 24 presents the gameplay scores on Pokémon Red. Gemini-2.5-pro achieves the highest score of 83.3, followed by Deepseek-R1 (75.0) and Claude-3.7 (63.9). GPT-4o achieves a score of 38.9. Notably, all the open-source LLMs evaluated (Llama-3.2-1B/3B, Qwen-2.5-3B/7B, and Minitron-4B/8B) recorded a score of 0.0, indicating significant challenges in playing Pokémon Red. Interestingly, o3-mini, which is expected to perform well due to its reasoning capabilities, also achieved a score of 0.0. We observed that o3-mini exhibited a tendency to rely on its pre-existing knowledge or intuition rather than adapting to the game environment, such as consistently moving downwards based on a likely incorrect assumption about the exit’s location, leading to unproductive repeated actions. This highlights the difficulty some models face in grounding their reasoning within the specific context of the game.
Table 24 shows the impact of reflection and planning modules. For GPT-4o, the ‘reflection-planning’ agent achieved the highest score (38.9), followed by the ‘reflection’ agent (36.1), and then the ‘planning’ and ‘zero-shot’ agents (both at 33.3). This suggests that reflection plays a crucial role in improving performance in Pokémon Red. In contrast, Llama-3.2-3B consistently scored 0.0 across all agent configurations.
The comparison across input modalities is presented in Table 24. For Gemini-2.5-pro and Claude-3.7, using both text and image inputs resulted in performance equal to or better than using text input alone. This indicates that visual information can be beneficial in this environment. GPT-4o also showed a performance increase when both modalities were used (41.7) compared to text-only input (38.9), suggesting that incorporating image data aids the agent’s decision-making.
| Models | Pokémon Red | Rank |
| Llama-3.2-1B | 0.00.0 | 8.5 |
| Llama-3.2-3B | 0.00.0 | 8.5 |
| Qwen-2.5-3B | 0.00.0 | 8.5 |
| Qwen-2.5-7B | 0.00.0 | 8.5 |
| Minitron-4B | 0.00.0 | 8.5 |
| Minitron-8B | 0.00.0 | 8.5 |
| GPT-4o-mini | 0.00.0 | 8.5 |
| GPT-4o | 38.99.6 | 4 |
| o3-mini | 0.00.0 | 8.5 |
| Gemini-2.5-pro | 83.30.0 | 1 |
| Claude-3.7 | 63.99.2 | 3 |
| Deepseek-R1 | 75.00.0 | 2 |
| Models | Agent | Pokémon Red | Rank |
| Llama-3B | Zero-shot | 0.00.0 | 6.5 |
| Reflection | 0.00.0 | 6.5 | |
| Planning | 0.00.0 | 6.5 | |
| Ref-Plan | 0.00.0 | 6.5 | |
| GPT-4o | Zero-shot | 33.30.0 | 3.5 |
| Reflection | 36.14.8 | 2 | |
| Planning | 33.30.0 | 3.5 | |
| Ref-Plan | 38.99.6 | 1 |
| Models | Input | Pokémon Red | Rank |
| GPT-4o | Text | 38.99.6 | 6 |
| Both | 41.78.3 | 5 | |
| Gemini | Text | 83.30.0 | 1.5 |
| Both | 83.30.0 | 1.5 | |
| Claude | Text | 63.919.2 | 4 |
| Both | 72.24.8 | 3 |
Appendix H Darkest Dungeon
H.1 Game Description for Darkest Dungeon
Environment.
Darkest Dungeon is a turn-based roguelike role-playing game where the player manages a roster of heroes as they explore procedurally generated dungeons filled with monsters, traps, and treasures (Red Hook Studios, 2016).
Each hero has unique abilities and a stress level that influences their behavior during combat and exploration.
Stress accumulates through continued exploration and battle, and heroes who reach high stress thresholds may develop afflictions that hinder or occasionally enhance their performance.
The game emphasizes tactical positioning, turn-order strategy, and long-term roster management.
Elements of randomness, such as attack accuracy, critical hits, and affliction outcomes, introduce uncertainty and require players to adapt their strategies dynamically.
For implementation, we build our environment on top of a rule-based bot (Kgleken, 2023), replacing its rule-based combat logic with the decisions made by LLM agents.
To access internal game states, we utilize the Darkest Dungeon Save Editor (Robojumper, 2023).
Since the game does not support complete control via keyboard input, we employ the Xbox 360 controller emulator to inject actions from the LLM agent.
We evaluate the agent’s performance during the first embarkation mission after the tutorial, which we designate as our default evaluation setting.
For consistency, we fix the party roster to include the Plague Doctor, Vestal, Highwayman, and Crusader, in that order, and equip the inventory with an additional 8 ‘Food’ and 8 ‘Torches’.
To ensure reproducibility, we provide a save file with this setup preconfigured.
During the mission, dungeon exploration is handled by rule-based logic, while all combat decisions are delegated to the LLM agent.
Observation-to-Text Conversion. We convert the internal game state of Darkest Dungeon into a structured textual description suitable for LLM input. The observation includes combat-relevant details such as the active hero’s stats, available skills, party composition, and enemy formation. For each hero, we format key attributes (e.g., HP, stress, position, status effects) along with skill availability and target constraints, using symbolic descriptors extracted from a parsed skill configuration file. Enemy information is similarly structured, including HP, rank, resistances, and threat indicators. The final text is composed of three parts: a detailed hero description with skill information, a party summary with basic stats, and an enemy formation breakdown.
Action Space. At each decision point, the agent chooses one of four action types: ‘Attack’, ‘Heal’, or ‘Swap’. For ‘Attack’ and ‘Heal’ actions, the agent specifies a skill slot index and a target index corresponding to a specific enemy or ally. ‘Swap’ actions require the agent to provide the current hero’s rank and a swap distance. To allow for complex plans (e.g., swapping then healing), the agent may output up to two such structured command lines in sequence.
H.2 Gameplay Prompt for Darkest Dungeon
Figure 17 shows the action inference prompt used by the ‘zero-shot’ agent for playing Darkest Dungeon. The system prompt encodes task-specific knowledge, including (1) the primary objective of completing expeditions while minimizing party stress, (2) strict targeting constraints for combat skills based on hero and enemy rank positions, and (3) the expected format for issuing valid commands. The user prompt provides the current game state, including the acting hero’s stats and available skills, a summary of the party, and the enemy formation. Some fields are omitted for brevity, but the format mirrors the actual prompt used during inference.
H.3 Evaluation Metric for Darkest Dungeon
To evaluate the performance of an LLM agent in Darkest Dungeon, we define a composite scoring metric that reflects progress, survivability, and stress management throughout the expedition. The score consists of three components: (1) the proportion of room combats successfully cleared, weighted at 40 points, (2) the fraction of heroes who survive the entire mission (out of four), weighted at 30 points, and (3) the remaining stress capacity of the team, also weighted at 30 points. However, the latter two components are only counted if the stage is successfully cleared (i.e., if the first term reaches its full 40 points); otherwise, they are set to zero. To ensure fairness, the stress of any hero who dies before the end of the run is treated as the maximum stress (200) when computing the average team stress. The final score is computed as:
H.4 Experimental Configuration for Darkest Dungeon
For all 6 open-source LLMs, we use a temperature of 0.7 and a repetition penalty of 1, and set the maximum number of game steps to 200. We run all experiments with 3 trials and report the average score with the standard deviation.
H.5 Result for Darkest Dungeon
| Models | DarkestD | Rank |
| Llama-3.2-1B | 0.00.0 | 11.5 |
| Llama-3.2-3B | 47.539.2 | 9 |
| Qwen-2.5-3B | 44.822.2 | 10 |
| Qwen-2.5-7B | 88.82.0 | 6 |
| Minitron-4B | 0.00 | 11.5 |
| Minitron-8B | 63.830.4 | 8 |
| GPT-4o-mini | 81.35.8 | 7 |
| GPT-4o | 93.41.5 | 2 |
| o3-mini | 89.02.1 | 5 |
| Gemini-2.5-pro | 93.71.6 | 1 |
| Claude-3.7 | 89.92.5 | 4 |
| Deepseek-R1 | 91.71.1 | 3 |
| Models | Agent | DarkestD | Rank |
| Llama-3B | Zero-shot | 47.539.2 | 7 |
| Reflection | 47.339.0 | 8 | |
| Planning | 56.323.6 | 6 | |
| Ref-Plan | 57.031.6 | 5 | |
| GPT-4o | Zero-shot | 93.41.5 | 1 |
| Reflection | 85.210.3 | 3 | |
| Planning | 82.08.6 | 4 | |
| Ref-Plan | 91.62.5 | 2 |
| Models | Input | DarkestD | Rank |
| GPT-4o | Text | 93.41.5 | 2 |
| Image | 92.23.0 | 3.5 | |
| Gemini | Text | 93.71.6 | 1 |
| Image | 92.21.8 | 3.5 | |
| Claude | Text | 89.92.5 | 6 |
| Image | 90.15.7 | 5 |
Table 27 reports the gameplay scores of various models on Darkest Dungeon. We show that Gemini-2.5-pro achieves the highest score of 93.7, closely followed by GPT-4o (93.4) and Deepseek-R1 (91.7), demonstrating strong capabilities across combat decisions and roster management. Among open-source models, Qwen-2.5-7B performs best, achieving a score of 88.8. In contrast, smaller models such as Llama-3.2-1B and Minitron-4B fail to make meaningful progress, often producing invalid outputs and scoring 0.0. A closer analysis reveals that Llama-3.2-1B frequently fails to follow the correct action format, while Llama-3.2-3B and Qwen-2.5-3B tend to issue invalid commands, such as using unavailable skills from incorrect hero positions or targeting unreachable enemies. These models also overuse ‘Swap’ actions, leading to inefficient combat sequences. Notably, many small models become stuck in a loop when the ‘Crusader’ hero is affected by the ‘Surprised!’ status and repositioned to the back row. Since most of the Crusader’s skills are unusable from that position, the models repeatedly attempt invalid actions or swap ineffectively, wasting turns and failing to recover from the disrupted formation.
Table 27 and Table 27 present ablation studies on agentic modules and input modalities in Darkest Dungeon. For agentic components, GPT-4o performs best in the zero-shot setting, with reflection and planning offering marginal or even negative impact on its performance. Llama-3.2-3B shows small but consistent gains when equipped with the planning module, though overall improvements remain limited. This suggests that large models like GPT-4o already possess sufficient planning capabilities for this task, while smaller models benefit only modestly from explicit agentic prompting. In terms of modality (Table 27), we find that providing image input in addition to text yields minimal improvement. For all models, performance remains similar or only slightly better when image data is included, indicating that the agents primarily rely on the structured text input to make decisions.
Appendix I Minecraft
I.1 Game Description for Minecraft
Environment. Minecraft (Mojang Studios, 2011) is an open-ended sandbox game where players explore a world, gather resources, and survive by placing and breaking blocks. This environment is based on the Mineflayer JavaScript API (contributors, 2013). Using Mineflayer, the agent can control a Minecraft bot through high-level JavaScript commands. We use Minecraft version 1.19 for compatibility, and to ensure consistent evaluation, we fix the world seed to 42 and initialize the bot at coordinates (604, 100, -823). The bot starts in survival mode with no items and progressively crafts a target item using its in-game observations and JavaScript-based actions generated by the LLM. We select 8 target items with varying levels of crafting difficulty: ‘crafting table’, ‘stone pickaxe’, ‘furnace’, ‘bucket’, ‘golden sword’, ‘diamond pickaxe’, ‘enchanting table’, and ‘nether portal’.
Observation-to-Text Conversion. The Mineflayer API provides the bot with its state and contextual information from the surrounding environment. Specifically, the bot receives textual observations in the form of: {Current biome, DayTime, Nearby blocks, Health status, Hunger status, Position, Equipped items, Inventory contents}.
Action Space. The action space consists of JavaScript code that interfaces with the Mineflayer API. Following Voyager (Wang et al., 2023), we expose the following set of control primitives to guide the LLM in generating valid and effective code actions for the bot using in-context learning.
exploreUntil(bot, direction, maxTime, callback): Moves the agent in a fixed direction for up to maxTime seconds, or until a custom stopping condition (defined in callback) is satisfied.
mineBlock(bot, name, count): Mines and collects up to count number of blocks with the specified name, within a 32-block radius.
craftItem(bot, name, count): Crafts the specified item using a nearby crafting table.
placeItem(bot, name, position): Places a block of the specified type at the given position.
smeltItem(bot, itemName, fuelName, count): Smelts the specified item using the provided fuel. Requires access to a nearby furnace.
KillMob(bot, mobName, timeout): Hunts and eliminates the specified mob within the time limit, and collects any resulting drops.
getItemFromChest(bot, chestPosition, itemsToGet): Navigates to the chest at the given location and retrieves the requested items.
depositItemIntoChest(bot, chestPosition, itemsToDeposit): Navigates to the given chest and deposits specified items into it.
I.2 Gameplay Prompt for Minecraft
Figure 18 shows the action inference prompt used by the ‘zero-shot’ agent for playing Minecraft. The system prompt contains gameplay-specific knowledge and guidance for action inference using Mineflayer APIs. It includes (1) the main task of the game, (2) control primitives, which is Javascript code template that should be referred to, (3) game observation in text format, and (4) the expected output response format with reasoning and code. The user prompt provides the current game state provided by Mineflayer APIs. Given this prompt, the LLM agent infers the appropriate Javascript code for action to complete the target task.
I.3 Evaluation Metric for Minecraft
We evaluate agent performance using the success rate of crafting target item. Since Mineflayer APIs provide access to the agent’s inventory, we compute the success score by checking whether the target item appears in the inventory at each game step. If the item is successfully crafted and presented, the episode is marked as successful with a score of 100; otherwise, 0.
I.4 Experimental Configuration for Minecraft
For all LLMs, we use a temperature of 1 and a repetition penalty of 1. The interaction with the game environment is limited to a maximum of 100 steps. We repeat all experiments 3 times to craft each target item, and report the average score and standard deviation.
I.5 Result for Minecraft
| Models | Minecraft | Rank |
| Llama-3.2-1B | 0.00.0 | 9.5 |
| Llama-3.2-3B | 0.00.0 | 9.5 |
| Qwen-2.5-3B | 0.00.0 | 9.5 |
| Qwen-2.5-7B | 0.00.0 | 9.5 |
| Minitron-4B | 0.00.0 | 9.5 |
| Minitron-8B | 0.00.0 | 9.5 |
| GPT-4o-mini | 46.07.0 | 5 |
| GPT-4o | 71.07.0 | 4 |
| o3-mini | 75.00.0 | 2 |
| Gemini-2.5-pro | 75.00.0 | 2 |
| Claude-3.7 | 75.00.0 | 2 |
| Deepseek-R1 | 41.70.0 | 6 |
| Models | Agent | Minecraft | Rank |
| Llama-3B | Zero-shot | 0.00.0 | 6 |
| Reflection | 0.00.0 | 6 | |
| Planning | 0.00.0 | 6 | |
| Ref-Plan | 0.00.0 | 6 | |
| GPT-4o | Zero-shot | 0.00.0 | 6 |
| Reflection | 50.00.0 | 1.5 | |
| Planning | 13.00.0 | 3 | |
| Ref-Plan | 50.00.0 | 1.5 |
As shown in Table 29, o3-mini, Gemini-2.5-pro, and Claude-3.7-sonnet performed the best on Minecraft, each obtaining a score of 75.0. All three models successfully crafted the following six items: ‘crafting table’, ‘stone pickaxe’, ‘furnace’, ‘bucket’, ‘golden sword’, and ‘diamond pickaxe’. However, they all failed to craft the two most difficult items: ‘enchanting table’ and ‘nether portal’. In contrast, all six open-source LLMs failed to craft any of the target items, resulting in a score of 0.0. Notably, five of the models, except Minitron-8B, failed to generate any executable JavaScript code compatible with the Mineflayer API. While Minitron-8B was able to generate valid code sometimes to move the bot and mine wood, it failed to craft even the simplest item, the crafting table.
As shown in Table 29, the reflection module, which encourages the model to generate improved code actions based on past failed attempts, significantly improved the performance of GPT-4o. However, the ‘reflection-planning’ agent achieved a score of 50.0, which is lower than the default ‘skill-management’ agent score of 71.0 in Table 29. This suggests that the skill management module, which is responsible for storing previously successful code actions and retrieving them when needed, plays a more substantial role in enhancing the performance of Minecraft.
Appendix J Stardew Valley
J.1 Game Description for Stardew Valley
Environment. Stardew Valley (ConcernedApe, 2016) is a life simulation and farming role-playing game. The player can engage in a variety of daily activities such as farming, fishing, mining, foraging, and socializing with villagers.
Our objective is to evaluate an LLM agent’s ability to autonomously perform farming-related tasks that maximize monetary gain within the first 13 in-game days (i.e., until the Egg Festival on Spring 13). Specifically, we focus on harvesting crops and strategically earning money by predicting high-profit crops and interacting with the in-game environment. The character begins on Day 1 with 200 gold. Four types of seeds are available for purchase: parsnip seeds, bean starter, cauliflower seeds, and potato seeds. Each seed type differs in cost, selling price, days to harvest, and other characteristics. Table 30 summarizes the properties of each seeds. To make the task challenging and enforce planning under resource constraints, we manually set the player’s maximum energy to 50 (default: 200). Using tools such as the hoe or watering can consumes 2 energy per use, limiting the number of tiles the agent can till or water in a single day. If the agent’s energy drops to 0 or below, they start the next day with only 26 energy (instead of 50). Furthermore, if energy falls -15 or below, the player loses 10% of their current gold and still begins the following day with only 26 energy. This constraint encourages the agent to prioritize actions and manage resources efficiently.
| Seed | Buy Price | Sell Price | Growth Days | Notes |
| Parsnip Seeds | 20 | 35 | 4 | |
| Bean Starter | 60 | 40 | 10 | Regrows every 3 days after first harvest |
| Cauliflower Seeds | 80 | 175 | 12 | |
| Potato Seeds | 50 | 80 | 6 | 20% chance of extra yield |
We run the game on the Steam platform and use the modding tool SMAPI (Pathoschild, 2025) to extract in-game states and implement custom actions. To send keyboard and mouse inputs, we use the pyautogui library on macOS and the AutoHotkey (AHK) library on Windows, following prior work (Tan et al., 2024). Since many in-game actions such as planting or watering crops consist of multiple low-level actions (e.g., move up/down/right/left, switch tool, use tool), we define a set of high-level actions to abstract these into semantically meaningful units. Each high-level action is mapped to a predefined sequence of keyboard inputs and implemented via SMAPI scripts with custom keyboard bindings. We provide the save point used for our experiment to ensure reproducibility.
Observation-to-Text Conversion. After each high-level action is executed, we extract a JSON-formatted game state via SMAPI, which includes information such as player location, current inventory, crop states in the field, remaining energy, and money. This state is then serialized into a natural language description and passed to the LLM.
Action Space. We define a compact action space consisting of 8 high-level actions essential for solving the task. These actions abstract away low-level controls and are defined as follows:
till_soil(num_tiles): Tills num_tiles soil tiles to prepare them for planting. The tiles are selected in a fixed order starting from the pre-defined position.
plant_seeds(): Plants all available seeds from the inventory into empty, tilled soil tiles. If the number of tilled tiles is less than the number of seeds, only the available plots are used.
water_seeds(): Waters all planted crops that have not yet been watered on the current day.
harvest_crops(): Harvests all crops that are fully grown and ready to be collected.
sell_item(): Sells all harvested crops currently in the inventory.
buy_item(item_name, item_count): Opens the shop interface, selects the specified item, and attempts to purchase the specified quantity. If there is insufficient money, the agent buys as many units as possible.
get_out_of_house(): Moves the character out of the house.
go_house_and_sleep(): Navigates the character back to the house, enters it, moves to the bed, and interacts with it to end the day.
J.2 Gameplay Prompt for Stardew Valley
Figure 19 shows the action inference prompt used by the ‘zero-shot’ agent for playing Stardew Valley. The system prompt defines the agent’s role as an in-game assistant tasked with selecting the best next action based on the current situation and target task. It specifies strict behavioral rules, such as only using actions from a predefined set, avoiding repeated failed actions, and formatting outputs as Python code (up to two actions). The valid action set includes available actions with clear descriptions and constraints. The user prompt provides the goal and detailed context, including crop stats, energy rules, and the current game state (location, inventory, weather, etc.). It also includes the last executed actions and requires the agent to valid action output.
J.3 Evaluation Metric for Stardew Valley
The objective in the Stardew Valley task is to maximize the amount of money earned during the first 13 in-game days, starting from Day 1. This period is a natural milestone in the game, as the Egg Festival takes place on Spring 13, where players can purchase high-reward crops such as Strawberry seeds. We evaluate performance based on the net profit earned by the end of Spring 13, calculated as the difference between the final gold amount and the initial amount of 200 gold. We normalize the score using a human expert baseline of gold, which we assign a normalized score of 100. Formally, the normalized score is defined as
where is the agent’s final gold amount, is the starting gold, and is the human expert score.
J.4 Experimental Configuration for Stardew Valley
For all LLMs, we use a temperature of 0.0 and a repetition penalty of 1. The interaction with the game environment is limited to a maximum of 150 steps. For GPT-4o, we used GPT-4o-2024-05-13 model for Stardew Valley. During LLM inference, the game is paused to prevent in-game time from progressing, which could otherwise alter the environment (e.g., a day transition). We perform each experiments three times and and present the average score with the standard deviation.
J.5 Result for Stardew Valley
| Models | Stardew | Rank |
| Llama-3.2-1B | 0.00.0 | 10.5 |
| Llama-3.2-3B | 0.00.0 | 10.5 |
| Qwen-2.5-3B | 0.00.0 | 10.5 |
| Qwen-2.5-7B | 0.00.0 | 10.5 |
| Minitron-4B | 0.00.0 | 10.5 |
| Minitron-8B | 0.00.0 | 10.5 |
| GPT-4o-mini | 16.127.8 | 7 |
| GPT-4o | 81.44.8 | 2 |
| GPT-5 | 92.38.6 | 1 |
| o3-mini | 55.116.0 | 5 |
| Gemini-2.5-pro | 59.210.1 | 4 |
| Claude-3.7 | 53.620.9 | 6 |
| Deepseek-R1 | 66.111.5 | 3 |
| Models | Agent | Stardew | Rank |
| Llama-3B | Zero-shot | 0.00.0 | 6.5 |
| Reflection | 0.00.0 | 6.5 | |
| Planning | 0.00.0 | 6.5 | |
| Ref-Plan | 0.00.0 | 6.5 | |
| GPT-4o | Zero-shot | 40.535.2 | 3 |
| Reflection | 18.35.2 | 4 | |
| Planning | 64.623.7 | 2 | |
| Ref-Plan | 95.75.7 | 1 |
| Models | Input | Stardew | Rank |
| GPT-4o | Text | 81.44.8 | 1 |
| Image | 0.00.0 | 8.5 | |
| Both | 41.922.1 | 6 | |
| Gemini | Text | 59.210.1 | 3 |
| Image | 7.69.0 | 7 | |
| Both | 60.06.0 | 2 | |
| Claude | Text | 53.620.9 | 4 |
| Image | 0.00.0 | 8.5 | |
| Both | 49.81.0 | 5 |
Table 33 presents a comparison of LLM performance in Stardew Valley. Among the models, GPT-4o presents the best performance, followed by Gemini-2.5-pro. Open-sourced LLMs fails to earn money, primarily for two reasons; (1) failure to perform valid actions (all models excepts Qwen-2.5-7B) (2) poor crop scheduling, resulting in crops not being ready for harvest on Day 13 (Qwen-2.5-7B). Under the imposed energy constraints, the optimal strategy is to plant Parsnip Seeds every four days, as they yield the highest profit. The player should purchase and plant as many seeds as possible on Days 1 and 5, and exactly 24 seeds on Day 9. Planting more than 24 seeds on Day 9 depletes the player’s energy during watering, leading to insufficient energy the next day and ultimately a failure to harvest on Day 13. None of the LLMs, including the API-based models, fully followed this optimal strategy. In particular, most models frequently selected suboptimal crops such as potato seeds, which contributed to their lower performance.
Table 33 shows the effect of different agentic modules on performance. In the case of GPT-4o, the planning module has a significant effect, as the task requires accurate seed selection and scheduling. In contrast, Llama-3.2-3B model fails to make a profit across four agent configuration, indicating that the underlying model’s capabilities are a limiting factor regardless of agent design.
Table 33 summarizes game performance under different input modalities. Notably, all three proprietary LLMs fails to achieve strong performance when only vision input is available. This underperformance is mainly due to the difficulty of extracting structured information from a single screenshot (see figure 19 for an example). Although the screenshot contains rich contextual information, it also contains a lot of redundant content, and critical information occupies only a small portion of the image. For instance, watered plants appears slightly darker than dry soil, making it difficult to distinguish visually. Similarly, the current day, a crucial cue for planning, is indicated in a small font in the upper-right corner of screenshot. When both text and image inputs are given, GPT-4o and Claude-3.7-Sonnet exhibit a performance drop, while Gemini-2.5-pro shows improved performance. This suggests that only sufficiently capable models can effectively integrate multimodal information, while others may struggle with modality fusion or become distracted by noisy visual inputs.
Appendix K StarCraft II
K.1 Game Description for StarCraft II
Environment. StarCraft II is a real-time strategy game where players gather resources, construct buildings, train units, and command armies to defeat opponents. The environment features a partially observable map, requiring the agent to explore and gather information about the opponent’s actions. For implementation, we adopt the BurnySc2/python-sc2 environment (BurnySc2, 2017), a Python interface widely used in the reinforcement learning (RL) community. Note that the library supports the raw scripted interface without a graphics-based interface. The environment supports various official maps and game modes. For our default evaluation (as in Table 3), we use the ‘Ancient Cistern LE’ map with the agent playing as Protoss against the built-in AI bot (Zerg, Hard difficulty; employing a timing build order strategy). The environment allows testing across different races, maps, and difficulty settings. For instance, the ‘Babylon LE’ is used to assess intra-game generalization in Table 7.
Observation-to-Text Conversion. The BurnySc2/python-sc2 environment provides the agent with observations capturing the current game state and context. Specifically, the observations include: {Resources, Buildings, Units, Research Progress, In-progress Actions, Enemy Information, Game Time}. We convert these observations into a concise text summary; an example summary is shown in Figure 20.
Action Space. Following the BurnySc2/python-sc2 implementation, We define the action space as a discrete set of 72 high-level commands specifically for the Protoss race. These include unit training (e.g., Probes, Zealots, and Stalkers), building construction (e.g., Pylons and Gateways), research upgrades, scouting, multi-unit attacks or retreats, and special abilities (e.g., Chrono Boost). A complete list of these commands is provided in Figure 20. At each game step, the agent generates a list of five actions, which are executed in order.
K.2 Gameplay Prompt for StarCraft II
Figure 20 shows the action inference prompt used by the ‘zero-shot’ agent for playing StarCraft II. The system prompt contains gameplay-specific knowledge and detailed instructions for action inference tailored to the Protoss race. It includes (1) the main task and game context, (2) a comprehensive action dictionary listing all possible unit production, building construction, and research actions, (3) the current game status summary including resources, buildings, units, and ongoing actions, and (4) the expected output response format that guides the agent to provide a step-by-step analysis, reasoning, and 5 concrete actionable commands with cost and resource availability considerations.
The user prompt provides the current game state with detailed information on resources, supply, buildings, units, and ongoing unit production. Given this prompt, the agent infers the appropriate Protoss-specific actions to optimize the gameplay strategy against a Zerg opponent, focusing on resource management, army composition, and tech progression to counter the enemy effectively.
K.3 Evaluation Metric for StarCraft II
Single-Agent Play. In the single-player mode, the agent competes in a series of matches against the AI bot opponent. It plays up to 4 matches, continuing until it either wins or loses. The evaluation metric is the win rate, calculated as:
Multi-Agent Play. In the multi-player mode, we do not use win rate as the performance metric, since only a single run is conducted per match. Instead, we measure the performance by the army supply difference between agents at the end of the match. The difference is calculated as the sum of each unit’s count multiplied by its consumed resource cost, reflecting the effective army strength. The winning agent receives a positive score equal to this army supply difference, while the losing agent is assigned the negative of this value. This scoring method captures not only victory but also the margin of the win.
Using these difference-based scores, we then compute Elo ratings for all agents following the Bradley-Terry model (Bradley & Terry, 1952), following the approach used in StreetFighter III multi-agent evaluation in Section C.3.
The resulting army supply difference matrix and Elo scores are presented in Figure 4(b).
K.4 Experimental Configuration for StarCraft II
For all LLMs, we use a temperature of 0.1 and a repetition penalty of 1.0. During LLM inference, we pause the StarCraft II game environment. Interactions with the game environment are limited to a maximum of 1,000 steps. For single-agent play, we repeat the experiments 4 times and report the average score along with the standard deviation. For multi-agent play, we run a single experiment and report the Elo score.
K.5 Result for StarCraft II
Single-Agent Play. We evaluate agent performances in a single-agent environment, all operating in a ref-plan setting. Table 36 shows a significant performance gap between open-source LLMs and proprietary LLMs. Among them, Llama-3.2-1B/3B models mostly repeat simple actions like scouting and mining minerals, without showing much strategic planning. On the other hand, proprietary LLMs, except for o3-mini, achieve over 50% win rate. Notably, GPT-4o and Gemini-2.5-pro won all four matches against the AI bot. They demonstrate strategic behavior by appropriately allocating resources over time in line with the game’s progression.
Table 36 presents ablation studies on agentic modules. As previously mentioned, Llama-3.2-3B models fail to manage matches effectively regardless of ablation settings, resulting in a 0% win rate. In contrast, GPT-4o demonstrates remarkable planning abilities, which are critical in StarCraft II given its real-time strategy nature. In other words, long-term planning to sustain strategies over time plays a pivotal role in securing victories. Interestingly, models relying solely on reflection perform worse than zero-shot, suggesting that reflection without proper planning may actually degrade performance.
Table 36 presents the results of modality experiments. Surprisingly, for GPT-4o, using both text and image inputs results in decreased performance, and Gemini and Claude also show no improvement. This implies that agents can make sufficiently accurate decisions based on textual observations alone, and the addition of image input may introduce challenges in multimodal reasoning.
Multi-Agent Play. We evaluate agent performances in a multi-agent setting by conducting pairwise matches among seven LLMs (excluding Gemini-2.5-pro), all operating under a ref-plan configuration. The results are summarized in Figure 4(b). To ensure a fair comparison, both competing agents consistently use the Protoss race in every match.
Interestingly, unlike the single-agent evaluation where Claude achieved only a 50% win rate (Table 36), Claude outperforms GPT-4o and attains the highest Elo rating in the multi-agent arena. Even more surprising is that Minitron-8b, which had 0% win rate in single-agent play, defeated GPT-4o, o3-mini, and GPT-4o-mini, earning the second highest Elo rating. This discrepancy suggests that the presence of multiple intelligent agents can significantly alter game dynamics, potentially due to increased strategic diversity or emergent adversarial behaviors. We also acknowledge that a single evaluation episode may have introduced some bias in the observed rankings.
| Models | StarCraft II | Rank |
| Llama-3.2-1B | 0.00.0 | 9.5 |
| Llama-3.2-3B | 0.00.0 | 9.5 |
| Qwen-2.5-3B | 0.00.0 | 9.5 |
| Qwen-2.5-7B | 0.00.0 | 9.5 |
| Minitron-4B | 0.00.0 | 9.5 |
| Minitron-8B | 0.00.0 | 9.5 |
| GPT-4o-mini | 75.050.0 | 3 |
| GPT-4o | 100.00.0 | 1.5 |
| o3-mini | 25.050.0 | 6 |
| Gemini-2.5-pro | 100.00.0 | 1.5 |
| Claude-3.7 | 50.057.7 | 4.5 |
| Deepseek-R1 | 50.057.7 | 4.5 |
| Models | Agent | StarCraft II | Rank |
| Llama-3B | Zero-shot | 0.00.0 | 6 |
| Reflection | 0.00.0 | 6 | |
| Planning | 0.00.0 | 6 | |
| Ref-Plan | 0.00.0 | 6 | |
| GPT-4o | Zero-shot | 50.057.7 | 2.5 |
| Reflection | 0.00.0 | 6 | |
| Planning | 50.057.7 | 2.5 | |
| Ref-Plan | 100.00.0 | 1 |
| Models | Input | StarCraft II | Rank |
| GPT-4o | Text | 100.00.0 | 2 |
| Both | 50.057.7 | 5 | |
| Gemini | Text | 100.00.0 | 2 |
| Both | 100.00.0 | 2 | |
| Claude | Text | 50.057.7 | 5 |
| Both | 50.057.7 | 5 |
Appendix L Slay the Spire
L.1 Game Description for Slay the Spire
Environment. Slay the Spire is a deck-building rogue-like game where the player ascends a procedurally generated three-act tower. Each act consists of a branching map with various room types such as combat encounters, shops, treasure rooms, rest sites, and random events, ending in a boss fight.
Our goal is to evaluate an LLM agent’s ability to reason over strategic choices in a stochastic, multi-step environment. Specifically, we task the agent with playing as the Ironclad character under standard rules (no ascension levels) and aim to defeat the final boss at floor 50. The LLM agent is responsible for two key decision types: (1) choosing which cards to play during combat, and (2) selecting rewards after battles. All other game decisions, such as map navigation, non-combat events, and potion usages, are handled by a simple rule-based policy.
The game runs on the Steam platform, and we use modding tools to enable communication between the game and an external agent. In particular, we use BaseMod (Bug Kiooeht, 2018a) and ModTheSpire (Bug Kiooeht, 2018b), along with a modified version of CommunicationMod (Forgotten Arbiter, 2019), which enables state extraction and action input via standard input/output streams. Our modified version also extracts detailed in-game information, including full card and relic descriptions.
To ensure consistency and reproducibility, we fix the game seed for all runs. This guarantees the same map layout, encounters, and card offerings. Additionally, since the original game unlocks card pools progressively as the player completes runs, we pre-unlock all cards using the Unlock Everything mod to make the full pool available from the start.
Observation-to-Text Conversion. After each action, we extract a JSON representation of the current game state, including player status, opponent status, current cards in hand, and relics. These states are serialized into a concise natural language summary that is passed to the LLM.
Action Space. Our environment defines two categories of actions: combat and card selection. During combat, the agent can either play a card (PLAY) or end its turn (END). In the card selection stage, the agent can choose a card reward (CHOOSE) or skip the reward (SKIP).
PLAY CARD_INDEX: Play a non-target card from the hand at position CARD_INDEX.
PLAY CARD_INDEX TARGET_INDEX: Play a targeted card from the hand at position CARD_INDEX, targeting opponent at TARGET_INDEX.
END: End the current turn.
CHOOSE CARD_INDEX: Select the card reward at position CARD_INDEX.
SKIP: Skip the card reward.
L.2 Gameplay Prompt for Slay the Spire
Figure 21 presents the action inference prompt utilized by the ‘zero-shot’ agent to play Slay the Spire. The system prompt describes the agent’s role as a strategic player in the game and outlines key game mechanics, including block, energy, card draw, and enemy intents. Detailed game states, such as player status, relic, card, and opponents are provided in the user prompt.
L.3 Evaluation Metric for Slay the Spire
The primary objective is to reach and defeat the final boss located on floor 50. Accordingly, the baseline score is determined by the highest floor cleared. For example, if the character dies on floor 43, the score would be 42. Since each floor has varying difficulty levels, and bosses—appearing at the end of each act—pose significant challenges, we assign bonus points to boss defeats. There are three major bosses and we grant an additional points for each boss defeated. Formally, the total score is defined as:
L.4 Experimental Configuration for Slay the Spire
For all LLMs, we use a temperature of 0.0 and a repetition penalty of 1. Interaction with the game environment is limited to a maximum of 200 steps. The game is not paused during LLM inference, as the game state does not change over time. All experiments carried out in three runs.
L.5 Result for Slay the Spire
| Models | SlaySpire | Rank |
| Llama-3.2-1B | 0.00.0 | 10 |
| Llama-3.2-3B | 0.00.0 | 10 |
| Qwen-2.5-3B | 0.00.0 | 10 |
| Qwen-2.5-7B | 5.00.0 | 6 |
| Minitron-4B | 0.00.0 | 10 |
| Minitron-8B | 0.00.0 | 10 |
| GPT-4o-mini | 3.32.9 | 7 |
| GPT-4o | 23.622.1 | 3 |
| o3-mini | 15.00.0 | 4.5 |
| Gemini-2.5-pro | 51.931.9 | 1 |
| Claude-3.7 | 15.00.0 | 4.5 |
| Deepseek-R1 | 24.917.1 | 2 |
| Models | Agent | SlaySpire | Rank |
| Llama-3B | Zero-shot | 0.00.0 | 6.5 |
| Reflection | 0.00.0 | 6.5 | |
| Planning | 0.00.0 | 6.5 | |
| Ref-Plan | 0.00.0 | 6.5 | |
| GPT-4o | Zero-shot | 12.34.6 | 4 |
| Reflection | 26.219.4 | 1 | |
| Planning | 23.214.2 | 3 | |
| Ref-Plan | 23.622.1 | 2 |
| Models | Input | SlaySpire | Rank |
| GPT-4o | Text | 23.622.1 | 3.5 |
| Both | 23.622.1 | 3.5 | |
| Gemini | Text | 51.931.9 | 1 |
| Both | 26.219.4 | 2 | |
| Claude | Text | 15.00.0 | 5 |
| Both | 9.74.6 | 6 |
Table 39 shows a comparison in gameplay performance between open-source LLMs and proprietary LLMs in Slay the Spire. Most open-sourced models fail to make meaningful progress in the game, resulting in near-zero scores. In contrast, API-based LLMs demonstrate significantly stronger performance: Gemini-2.5-pro achieves the highest score, defeating the second boss in two out of three runs. The inherent stochasticity of the game contributes to high variance across all proprietary models.
Table 39 presents an ablation study evaluating the impact of agentic modules, reflection and planning, on performance in Slay the Spire. For the weaker model (Llama-3.2-3B), none of the agent variants achieve any meaningful progress, indicating that architectural changes alone are insufficient without strong base capabilities. In contrast, GPT-4o benefits substantially from added reasoning modules: the reflection variant achieves the highest score, while both planning and reflection-planning agents also outperform the zero-shot baseline. Notably, the planning module is relatively less effective, which may stem from the fact that optimal actions are highly sensitive to the opponent’s intent, information that is only partially observable and difficult to predict multiple turns ahead. These results highlight that while base model capacity is a prerequisite, structured reasoning routines further enhance gameplay performance in complex decision-making environments.
Table 39 compares model performance in Slay the Spire when using either text-only input or a combination of text and vision inputs. In all three proprietary models, adding visual input does not improve performance—and in fact, often leads to degradation. This outcome is not entirely surprising, as crucial gameplay information, such as detailed card effects, relic descriptions, and power mechanics, is often absent in the game screenshot. As a result, the image input fails to provide meaningful utility and instead introduces ambiguity or redundancy, effectively acting as noise rather than useful context. These findings suggest that for structured, information-dense environments like Slay the Spire, high-quality textual representations remain the most reliable modality for LLM agents.
Appendix M Baba Is You
M.1 Game Description for Baba Is You
Environment. Baba Is You is a puzzle game in which players must discover and understand every rule and mechanic on their own, apart from the basic movement keys (‘left’, ‘right’, ‘up’, and ‘down’) (Hempuli, 2019). The game’s defining feature is that the text tiles forming the rules can be pushed around, allowing the player to rewrite those rules on the fly. Every valid rule sentence must contain a verb (e.g., ‘Is’ and ‘Has’), and text tiles with a colored background (e.g., ‘You’, ‘Push’ and ‘Win’) cannot serve as subjects. A level is cleared when the object designated by ‘You’ touches the object designated by ‘Win’. For instance, with the rules ‘Baba Is You’ and ‘Flag Is Win’, the player wins as soon as Baba touches the flag. For implementation, we instrument the Steam edition with a lightweight Lua mod that, after every player move, dumps the full internal game state, i.e., the coordinates of every object, to a JSON file. The Lua modding script uses mod hook functions provided by the game developer, which are triggered at specific points in the game’s code. The agent outputs are delivered to the game via simulated key presses using the pyautogui library. We use Level 1 - Where do I go?, shown in Figure 22, as our default evaluation setting (for Table 3).
Observation-to-Text Conversion. We present the LLM with a textual description of the game state in two parts. First, we list the current coordinates of every object in the level. Second, we manually parse the state to extract all active rules and append those rules to the prompt. The combined description forms the model’s observation.
Action Space. In the 2D grid environment, the agent can move ‘left’, ‘right’, ‘up’, or ‘down’. At each step, the agent outputs a finite sequence of these directions, which we translate into consecutive key presses and send to the game.
M.2 Gameplay Prompt for Baba Is You
Figure 23 shows the action inference prompt used by the ‘zero-shot’ agent for playing Baba Is You. The system prompt encodes key game-specific knowledge, including (1) the primary objective of touching a ‘Win’ object with a controllable (‘You’) object, (2) an explanation of how rules are formed and manipulated via text block arrangements, (3) crucial pushing mechanics with coordinate-based guidance to avoid unintended interactions, and (4) the expected input-output format for the LLM. The user prompt provides the current puzzle state, including the map dimensions, object locations, and active rules. Given this prompt, the LLM agent outputs a sequence of directional actions.
M.3 Evaluation Metric for Baba Is You
To evaluate the performance of an LLM agent, we define a hierarchical scoring metric that rewards meaningful progress toward solving the puzzle. As shown in Figure 22, a key prerequisite is breaking the rule ‘Wall Is Stop’, which enables movement out of the closed area. The second subgoal is creating a winning condition, such as forming the rule ‘Flag Is Win’, but this is only possible once the wall constraint is removed. Each subgoal provides 20 points. If the agent clears the level by having the ‘You’ object touch a ‘Win’ object, it receives a full score of 100, overriding subgoal rewards. The final score is computed as:
M.4 Experimental Configuration for Baba Is You
For all 6 open-source LLMs, we use a temperature of 0.7 and a repetition penalty of 1, and set the maximum number of game steps to 30. We run all experiments with 3 trials and report the average score with the standard deviation.
M.5 Result for Baba Is You
| Models | BabaIsYou | Rank |
| Llama-3.2-1B | 6.711.5 | 12 |
| Llama-3.2-3B | 20.00.0 | 6.5 |
| Qwen-2.5-3B | 13.311.5 | 10.5 |
| Qwen-2.5-7B | 20.00.0 | 6.5 |
| Minitron-4B | 20.00.0 | 6.5 |
| Minitron-8B | 20.00.0 | 6.5 |
| GPT-4o-mini | 13.311.5 | 10.5 |
| GPT-4o | 20.00.0 | 6.5 |
| o3-mini | 73.346.2 | 1.5 |
| Gemini-2.5-pro | 73.346.2 | 1.5 |
| Claude-3.7 | 46.746.2 | 3 |
| Deepseek-R1 | 20.00.0 | 6.5 |
| Models | Agent | BabaIsYou | Rank |
| Llama-3B | Zero-shot | 20.00.0 | 4.5 |
| Reflection | 20.00.0 | 4.5 | |
| Planning | 20.00.0 | 4.5 | |
| Ref-Plan | 20.00.0 | 4.5 | |
| GPT-4o | Zero-shot | 20.00.0 | 4.5 |
| Reflection | 20.00.0 | 4.5 | |
| Planning | 20.00.0 | 4.5 | |
| Ref-Plan | 20.00.0 | 4.5 |
| Models | Input | BabaIsYou | Rank |
| GPT-4o | Text | 20.00.0 | 6 |
| Image | 6.713.7 | 9 | |
| Both | 20.00.0 | 6 | |
| Gemini | Text | 73.346.2 | 2 |
| Image | 20.00.0 | 6 | |
| Both | 86.723.1 | 1 | |
| Claude | Text | 46.746.2 | 3 |
| Image | 20.00.0 | 6 | |
| Both | 20.00.0 | 6 |
The performance of different models on Level 1 of Baba Is You using the ‘reflection-planning’ agent is shown in Table 42. We observe that o3-mini and Gemini-2.5-pro achieve the highest score of 73.3, significantly outperforming all other models. Apart from these two reasoning models, only Claude-3.7-sonnet, a hybrid reasoning model, scores above 20.0, indicating that it is the only other model capable of constructing a valid winning condition. In contrast, all non-reasoning models, including every open-source small language model we tested, typically only managed to break the ‘Wall Is Stop’ rule, consistently earning a score of 20.0. However, a qualitative analysis of the models’ self-defined subtasks and reasoning traces suggests that these models often fail to infer that breaking ‘Wall Is Stop’ is a necessary prerequisite for constructing the winning condition. This implies that their success in breaking the rule was largely unintentional.
Table 42 presents an ablation study evaluating the impact of agentic modules on performance in Baba Is You. Across both Llama-3.2-3B and GPT-4o, we observe no measurable improvement over the zero-shot baseline, with all configurations achieving the same score of 20.0. This indicates that the agentic components do not significantly enhance the agent’s ability to reach the winning condition for these two models. The task remains challenging, largely due to the model’s limited spatial reasoning capabilities. While the models occasionally produce valid high-level plans, such as to form the rule ‘Flag Is Win’ at specific coordinates, they frequently fail to account for the game’s pushing mechanics. As a result, they often push text tiles in unintended directions, breaking or misaligning the intended rule formation.
Table 42 presents an ablation study on input modalities across several multimodal models. We observe that relying solely on image input significantly degrades performance for all models. Adding image input on top of text yields only marginal improvements, if any, suggesting that the agents primarily rely on text-based representations to make decisions. Notably, Gemini-2.5-pro benefits slightly from the combined input, achieving the highest score of 86.7.
Appendix N 2048
N.1 Game Description for 2048
Environment. 2048 (Cirulli, 2014) is a single-player sliding tile puzzle game played on a 4×4 grid. The objective is to combine numbered tiles by sliding them in one of four directions (i.e., up, down, left, or right) to create a tile with the value 2048. In this environment, the agent observes the current board state, represented as a 4×4 matrix of integers (each cell contains 0 for empty or a power of 2 for active tiles), and selects one of four discrete actions corresponding to directional moves. While the original game ends when no moves are available (i.e., the board is full and no adjacent tiles can be merged), we additionally terminate the episode if the agent performs five consecutive invalid moves (i.e., actions that result in no change to the board state). For implementation, we use an open-source, Pygame-based game environment. The logic module manages the board state, tile movements, merging, and win/loss conditions, while the interface leverages Pygame to render the board and handle user input. The implementation supports dynamic resizing and configurable parameters, and is designed to facilitate both human play and automated experiments.
Observation-to-Text Conversion. The environment’s board state can be directly transformed into a textual description, formatting it as a 4×4 array of integers in which each element indicates the value of the corresponding tile.
Action Space. The action space comprises four discrete actions: ‘up’, ‘down’, ‘left’, and ‘right’, each representing a possible direction in which the agent can slide the tiles on the board.
N.2 Gameplay Prompt for 2048
Figure 24 shows the action inference prompt and the corresponding game screenshot used by the zero-shot agent to play 2048. The system prompt includes (1) the main objective of the game, (2) detailed game rules, and (3) the expected input-output format between the LLM and the environment. The user prompt provides (1) the specific task for the 2048 game, (2) the previous board state and game score, (3) the last executed action, (4) the current board state and game score, and (5) the expected output format. Based on this information, the agent determines the next action to take.
N.3 Evaluation Metric for 2048
The goal of 2048 is to create a tile with the value 2048. The game score increases as tiles are merged, with the value of the merged tile added to the total score. Although the score when the 2048 tile is created can slightly vary depending on the gameplay, it is generally estimated that the score is around 20,000 points. Therefore, we define the evaluation metric as the progress to the target score of 20,000, normalized to 100. Formally, the normalized score is defined as:
where ‘Final Game Score’ denotes the total score at the end of the game. This metric reflects how close the agent came to achieving the primary objective of creating the 2048 tile.
N.4 Experimental Configuration for 2048
For all 6 open-source LLMs, including Llama-3.2-1B/3B, Qwen-2.5-3B/7B, and Minitron-4B/8B, we use a temperature of 0.0 and a repetition penalty of 1.0. We set the maximum number of game steps to 10,000. However, the maximum number of game steps (10,000) was never reached in the experiments. The game typically terminated either when the board was full and no adjacent tiles could be merged, or when the agent failed to take an action that changed the board state for more than five consecutive steps. In the best gameplay episode of o3 zero-shot agent, which achieved a score of 57.32, a total of 685 steps were taken to reach this result.
N.5 Result for 2048
All reported results in Tables 45–45 are computed as the mean and standard deviation over five independent runs, using a ‘zero-shot’ agent for each model configuration. To establish a baseline for comparison, we additionally included a random agent that selects one of the four possible actions (up, down, left, right) uniformly at random. This agent was evaluated over 50 episodes, and its performance is summarized in Tables.
| Models | 2048 | Rank |
| Random Agent | 5.52.3 | 6 |
| Llama-3.2-1B | 0.00.1 | 15 |
| Llama-3.2-3B | 0.30.2 | 12 |
| Qwen-2.5-3B | 0.10.1 | 13 |
| Qwen-2.5-7B | 0.60.4 | 11 |
| Minitron-4B | 0.10.0 | 14 |
| Minitron-8B | 0.70.7 | 10 |
| GPT-4o-mini | 1.11.0 | 9 |
| GPT-4o | 5.61.5 | 5 |
| o3-mini | 25.37.3 | 2 |
| o4-mini | 15.76.7 | 3 |
| o3 | 34.923.4 | 1 |
| Gemini-2.5-pro | 5.12.5 | 8 |
| Claude-3.7 | 5.32.7 | 7 |
| Deepseek-R1 | 11.53.4 | 4 |
| Models | Agent | 2048 | Rank |
| Random Agent | - | 5.52.3 | 6 |
| Llama-3B | Zero-shot | 0.30.2 | 8 |
| Reflection | 0.00.0 | 11 | |
| Planning | 0.10.1 | 10 | |
| Ref-Plan | 0.10.2 | 9 | |
| GPT-4o | Zero-shot | 5.61.5 | 5 |
| Reflection | 3.52.9 | 7 | |
| Planning | 6.05.5 | 4 | |
| Ref-Plan | 7.05.7 | 3 | |
| o3-mini | Zero-shot | 25.37.3 | 1 |
| Ref-Plan | 17.69.5 | 2 |
| Models | Input | 2048 | Rank |
| Random Agent | - | 5.52.3 | 5 |
| GPT-4o | Text | 5.61.5 | 3 |
| Image | 1.81.1 | 10 | |
| Both | 5.44.5 | 6 | |
| Gemini | Text | 5.12.5 | 8 |
| Image | 5.52.4 | 4 | |
| Both | 3.12.6 | 9 | |
| Claude | Text | 5.32.7 | 7 |
| Image | 8.44.0 | 1 | |
| Both | 6.70.9 | 2 |
Gameplay score on 2048. Beyond the 12 models presented in the main paper, we also include results for OpenAI’s more recent models, o4-mini and o3 in Table 45. Interestingly, none of the open-source models (e.g., Llama, Qwen, Minitron) were able to correctly interpret the 2D array prompt representing the game board. Consequently, they repeatedly issued the same action even when no tiles could be merged, leading to premature termination of the game.
In contrast, GPT-4o, Gemini-2.5-pro, and Claude-3.7 were able to detect invalid moves and avoid them to some extent. However, these models failed to manage merged tiles into their planning, often repeating superficially valid actions that quickly led to a board lock-up. As a result, their gameplay performance was almost indistinguishable from that of the random agent.
Notably, the reasoning-capable models—o3-mini, o4-mini, o3, and DeepSeek-R1—exhibited meaningful gameplay performance. While the open-source and non-reasoning models were typically limited to producing a maximum tile of 64 or 128, o3-mini successfully generated the 512 tile in 4 out of 5 runs, and o3 achieved the 1024 tile in 2 out of 5 runs.
A particularly interesting observation is that, even without explicit strategic instructions (e.g., cornering high-value tiles or arranging tiles in a staircase pattern), the reasoning-based models implicitly discovered human-like strategies based solely on the basic game rules provided via system prompts. This behavior is illustrated in Figure 25, where o3 exhibits an emergent form of spatial organization akin to that used by experienced human players.
Ablation study for agentic modules on 2048. In addition to the models evaluated in the main paper, we conducted an extended ablation study using the o3-mini model to assess the impact of agentic modules—Reflection and Planning—on gameplay performance in the 2048 environment. Table 45 presents the results of this study across three representative models: Llama-3.2-3B, GPT-4o, and o3-mini, under various agent configurations.
For Llama-3.2-3B and GPT-4o, enabling the Reflection or Planning modules—individually or together—did not yield significant gains, suggesting that external agentic scaffolding alone is insufficient for improving performance, likely due to limited reasoning or architectural constraints.
In contrast, o3-mini exhibited a counter-intuitive trend: the zero-shot agent significantly outperformed both the Reflection and the Reflection-Planning agents. A plausible explanation for this phenomenon is that reasoning-centric models such as o3-mini may already perform internal reflection and planning when generating a response. As such, the introduction of additional agentic modules may introduce redundancy or even propagate errors in intermediate reasoning steps, degrading performance.
Supporting this hypothesis, Figure 26 shows the inference output of o3 shortly before achieving a 1024-tile. The output reveals that the model implicitly forecasts and sequences the next three actions, essentially performing internal planning without explicit agent modules. This emergent behavior highlights the potential of foundation models with strong reasoning capabilities to handle multi-step decision problems autonomously without external agentic augmentation.
Comparison across modalities on 2048. To assess the impact of input modality on gameplay performance, we evaluated models using either textual or visual (image-based) representations of the 2048 board state. The results are summarized in Table 45. Interestingly, GPT-4o demonstrated weaker performance with image inputs compared to text, suggesting that its image understanding capabilities—at least within the structured context of 2048—may lag behind its text comprehension. This is consistent with prior findings that GPT-4o, while multimodal, exhibits varying levels of alignment across modalities depending on task complexity and structure. In contrast, both Gemini 2.5 Pro and Claude 3.7 achieved better performance with image-based inputs than with text. This indicates a stronger visual reasoning capability in these models, particularly when parsing structured 2D spatial layouts such as the 2048 board. Their ability to interpret and act on visual patterns appears to be more robust than their capacity to process raw 2D arrays expressed as textual input.These findings highlight that multimodal models exhibit non-uniform modality strengths, and task-specific evaluations are crucial for selecting the appropriate input format to maximize agent performance.
Appendix O Details for Data Augmentation
Prompting Details. To increase the diversity of our training dataset and avoid overfitting risks, we applied a data augmentation strategy focused on the system prompt. The original dataset was constructed by rolling out the same game prompt multiple times in the environment to collect assistant responses. The system prompt remains static, and the user prompt exhibits variations due to changes in game state, though these were constrained to a fixed format where only specific values changed.
To address this issue, we performed augmentation on the system prompt, which contains general information such as the LLM’s role, game rules, and behavioral guidelines. Since the user prompt was dependent on dynamic game states, modifying it risked introducing inconsistencies or hallucinations. Therefore, we chose to keep the user prompt fixed during augmentation.
We use GPT-4o to generate paraphrased versions of the original system prompt. Specifically, we prompted the model to produce 10 alternative phrasings of the original system prompt that preserved its semantics while varying its linguistic expression. Each paraphrased system prompt was then paired with the original user prompt and assistant response, resulting in 10 augmented versions of each original data point. Including the original, this expands the dataset by a factor of 11. Figure 27 presents an example of the original and augmented prompt in Slay the Spire.
This augmentation strategy helped increase the syntactic and lexical diversity of the dataset while preserving semantic fidelity and coherence, thereby supporting more robust fine-tuning of the LLM.
| of Augs. | SuperMario | 2048 |
| 0 | 24.2 | 1.4 |
| 3 | 26.8 | 3.5 |
| 10 | 26.7 | 2.8 |
| 20 | 25.9 | 2.4 |
Effect of The Number of Augmentations. Table 46 shows the fine-tuned model performance across different numbers of augmentations. We use the same fine-tuning setup as the intra-game generalization in Table 7. Overall, increasing the number of augmentations improves the performance of fine-tuned models compared to the one with no augmentation. Interestingly, the small number of augmentations (3 or 10) is quite effective to ensure linguistic diversity in fine-tuning, and the effectiveness diminishes as the number of augmentations grows.
Appendix P Implementation Details
P.1 Asynchronous Inferences for Multi-Agent Environment
Here we provide implementation details for multi-agent game environment that enables efficient, scalable, and realistic agent interaction.
Overview. The core design principle is to execute the game loop on a frame-by-frame basis while allowing each agent to asynchronously infer and initiate its next high-level action upon completion of the previous one. This approach is particularly well-suited for real-time, frame-based games such as Street Fighter III and StarCraft II, where individual actions may span a variable number of frames (e.g., a ‘Move Away’ action takes 4 frames, whereas a ‘Super Attack’ requires 7 frames to complete). In this system, each agent operates independently, without waiting for other agents to complete their actions. As soon as an agent completes its current action, it observes the current game state and determines its next high-level action. This asynchronous execution allows the game to progress fluidly and continuously, closely mimicking the dynamics of real-time multiplayer games.
Illustrative Example. Consider a multi-agent scenario in Street Fighter III. Agent 1 selects a ‘Super Attack’ that spans 7 frames, while Agent 2 chooses a ‘Move Away’ action that lasts 4 frames. The game orchestrator initiates both actions simultaneously. After 4 frames, Player 2’s action concludes, triggering the agent to observe the updated state and select a new action, e.g., a ‘Medium Punch’ lasting 2 frames. The orchestrator then integrates this new action into the ongoing simulation, even as Agent 1’s ”Super Attack” continues. This frame-by-frame orchestration proceeds iteratively. Each agent re-enters the decision-making process immediately upon completing its current action, independent of the progress or status of other agents.
Advantages. This asynchronous, non-blocking scheduling model provides several key advantages. It enables agents to operate with varying action durations without artificial synchronization barriers, facilitating a more natural and responsive interaction dynamic. The resulting system supports overlapping actions, better reflects the timing complexities of real-time games, and can be readily extended to support environments with more than two agents.
P.2 Fine-tuning
We conducted supervised fine-tuning using collected gameplay data to adapt the LLM agent to game-specific reasoning and interactions with environments.
| Task | LLaMA-3.2-1B(h) | LLaMA-3.2-3B(h) |
| Intra-game/Non-game | 1.7 | 4.2 |
| OOD-game | 1.7 | 4.3 |
Training Configuration. We fine-tune two models: Llama-3.2-1B and Llama-3.2-3B. Training is conducted using 4 NVIDIA A100 GPUs with 80GB of memory each. We use a learning rate of 1e-6, a per-device batch size of 4 with gradient accumulation steps set to 4, resulting in an effective batch size of 64. The models were trained for 1 epoch with 100 warm-up steps. Table 47 summarizes the GPU time taken for fine-tuning.
Data Statistics. In total, we use 105,502 data points for fine-tuning. Only data points containing fewer than 4,096 tokens are used for training to ensure compatibility with model input length limitations. All data points from Pokémon and 329 out of 9900 data points from Minecraft are discarded due to length constraint. For the out-of-distribution (OOD) game generalization experiments, we exclude the games 2048 and Super Mario, resulting in 87,660 data points used. Apart from the number of data points, the training configuration remained identical.
P.3 Unseen Scenarios
To evaluate the intra-game generalization capability of fine-tuned LLM agents, we define a separate scenario that is not used during the training dataset collection. Unseen scenarios differs from the seen ones by featuring a different character, map, or stage. Figure 28 presents example screenshots of the seen and unseen scenarios across six games.
Street Fighter III. While the character Ken was used during training data collection, a different character, Chun-Li, was introduced at evaluation time to assess intra-game generalization. Chun-Li features a distinct set of moves and hitboxes, posing a significantly different control and tactical challenge compared to Ken. The game environment and match settings remained unchanged.
Darkest Dungeon.
While the first expedition used during training data collection featured a party composed of a Plague Doctor, Vestal, Highwayman, and Crusader, the unseen scenario involved a different party composition: one Man-At-Arms, two Grave Robbers, and one Vestal. This change introduced significantly different combat dynamics, synergies, and positional requirements. Additionally, the enemy pool in this dungeon included the Madman—a monster not encountered during training—known for its stress-inducing attacks and erratic behavior. All other gameplay parameters, including the dungeon type (short dungeon in Ruins area), remained unchanged.
StarCraft II. During training, agents were exposed to the ‘Ancient Cistern LE’ map, while evaluation was conducted on a different map, ‘Babylon LE’ to assess intra-game generalization. All other game settings, including player race (Protoss), opponent race (Zerg), opponent’s build order strategy (Timing), and difficulty level (Hard), were kept constant.
Slay the Spire. We used game runs with different random seeds as unseen scenarios. Each game seed determines the layout of the map, including the sequence of opponents, bosses, events, and card reward options. All other game parameters, such as character choice (IronClad) and starting deck, were held constant.
Baba Is You. In the case of Baba Is You, the second stage, Level 2 - Now what is this?, is used as an unseen scenario. This stage features new rule combinations and object interactions that are absent from the training set, thereby testing the agent’s ability to generalize to novel logic structures. The game mechanics and control scheme remained unchanged.
Appendix Q Analysis of The Gap between In-game and OOD-games in Fine-tuning
| Genre | Action | Adventure | RPG | Simulation | Strategy | Puzzle |
| Games | SuperMario | AceAttorney | DarkestD | MineCraft | SlaySpire | 2048 |
| Llama-3.2-1B (Pretrain) | 18.7 | 1.3 | 0.0 | 0.0 | 0.0 | 0.1 |
| Llama-3.2-1B (Fine-tune) | 23.5 | 5.5 | 0.0 | 0.0 | 0.0 | 0.3 |
| Llama-3.2-3B (Pretrain) | 31.8 | 4.6 | 47.5 | 0.0 | 0.0 | 0.1 |
| Llama-3.2-3B (Fine-tune) | 24.9 | 9.1 | 40.1 | 0.0 | 0.0 | 3.1 |
| Avg. Improvement (%) | 2.0 | 210.5 | -7.8 | 0.0 | 0.0 | 1600.0 |
Setup. For studying the in-OOD gap in generalization, we fine-tune Llama-3.2 only on Baba Is You. Other hyperparameter configurations are exactly the same as in Appendix P.
Results. As shown in Table 48, fine-tuning yields the most improvements for OOD games in the same genre, but can also provide benefits across different genres in some cases. For example, 2048, another puzzle game, showed the largest improvement after fine-tuning. Interestingly, Ace Attorney, though a different genre, also benefited from fine-tuning, likely due to shared requirements in logical reasoning (see qualitative examples below). In contrast, for games like Darkest Dungeon and Super Mario, which involve different gameplay dynamics and action spaces, the effect was mixed. Note that, as shown in Table 7, unlike the current experiment (fine-tuned on Baba Is You only), LLMs fine‑tuned on 10 games (without Super Mario) showed consistent improvements on Super Mario (OOD game). This implies our fine‑tuning set, constructed from a diverse set of games across genres, contains knowledge broadly beneficial for gameplay/decision-making and supports cross‑genre generalizability.
Appendix R Limitations and Future Work
Cost Consideration. Among all games in Orak, six games, i.e., Ace Attorney, Her Story, Darkest Dungeon, Stardew Valley, Slay the Spire, and Baba is You, require a one-time purchase, typically priced ranging from to . While this represents a non-negligible upfront cost, it is relatively minor compared to the recurring cost associated with proprietary LLM API calls. From a cost-efficiency perspective, the benchmark remains accessible and practical for sustained research.
License Issue. We have made considerable efforts to comply with licensing requirements in designing our benchmark. (1) Users must purchase the commercial games themselves for evaluation. (2) We do not distribute any commercial game executables. (3) We do not modify any game assets during gameplay. Moreover, we explicitly state the following legal compliance guidelines “This project is strictly for research purposes. Researchers are required to use our framework solely as a benchmark for evaluating LLMs, and any models derived from our benchmark are strictly prohibited from commercial use”.
Real-time Gameplay. Street Fighter III, Super Mario, and StarCraft II inherently require real-time gaming, unlike other turn-based or simulation games in Orak. However, in our current evaluation setup, the game is paused during LLM inference to remove the impact of real-time constraints on agent performance. While this allows for more stable evaluation of reasoning capabilities, real-time responsiveness is critical in many gaming contexts, so it should be handled for practical needs. We leave latency-aware evaluation protocols for building real-time gaming LLM agents as future work.
Study on RL-based Fine-tuning. Although RL-based fine-tuning has demonstrated strong performance in many domains, such as mathematics and programming, we did not explore it in this study. Given the interactive nature of the Orak environments, it would be natural to derive dynamic, context-aware rewards from in-game feedback and apply RL fine-tuning methods such as DPO (Rafailov et al., 2023) or GRPO (Shao et al., 2024). Unlike domains such as mathematics or programming, where problems typically have static correct solutions derived through logical reasoning, gameplay requires strategic reasoning that adapts dynamically to the actions of other agents. In games, the optimal action is often contingent on the evolving behavior of the player or opponents, reflecting the complex, interactive nature of multi-agent environments. This distinction is particularly relevant to real-world domains such as business, economics, and negotiation, where strategic decision-making is frequently modeled using game-theoretic frameworks (Zhang et al., 2024). Therefore, leveraging Orak as an environment for RL-based fine-tuning may significantly enhance an LLM’s capacity to reason strategically and operate effectively in multi-agent settings. This line of research holds promise for improving LLM performance in a broad range of real-world applications where understanding multi-agent dynamics is essential.
Support for Diverse Modalities. Orak supports evaluation of LLMs and VLMs, but it does not extend to other modalities often essential in real-world gameplay. One key example is sound. For instance, in Minecraft, players rely on zombie audio cues to avoid danger, while in Street Fighter III and StarCraft II, alert sounds indicate some specific attacks by the enemy. Also, beyond games in Orak, First-Person Shooter (FPS) games usually use gunfire sounds for spatial awareness, and horror games often rely on audio to signal the proximity of threats. Consequently, incorporating other modalities such as audio remains an open challenge, and benchmarking the performance of emerging Speech-Language Models (Chu et al., 2023, Cui et al., 2024) in gaming scenarios could be an important step toward broader multimodal agent development.
Societal Impact. Simulation games, Minecraft and Stardew Valley, offer rich environments where players can explore, mine resources, and craft items, enabling life-like simulations of human behavior. These games offer a valuable testbed for analyzing long-horizon behavior of LLM agents by systematically comparing gameplay trajectories of humans and LLM agents, which enables a rigorous assessment of whether LLMs exhibit human-like decision-making patterns. Introducing multiple agents into these environments allows for the study of emergent social behaviors among LLM agents. We believe such settings are particularly well-suited for precisely measuring the social impact of complex agent behaviors, offering valuable insights into the dynamics of LLM-based agents.

### v3

@@
 We further release a fine-tuning dataset of expert LLM gameplay trajectories covering multiple genres, turning general LLMs into effective game agents.
-Orak offers a united evaluation framework, including game leaderboards, LLM battle arenas, and in-depth analyses of input modality, agentic strategies, and fine-tuning effects, establishing a foundation towards versatile gaming agents.
-Code is available at https://github.com/krafton-ai/Orak.
+Orak offers a united evaluation framework, including game leaderboards, LLM battle arenas, and ablation studies of input modality, agentic strategies, and fine-tuning effects, establishing a foundation towards versatile gaming agents.
+Code and datasets are available at https://github.com/krafton-ai/Orak and https://huggingface.co/datasets/KRAFTON/Orak.
 1 Introduction
-Large Language Model (LLM) agents are revolutionizing various industries (Wang et al., 2024a), and well-established benchmarks play a key role in unlocking their abilities on complex tasks, e.g., Coding (Hendrycks et al., 2021, Zhuo et al., 2024), Web Search (Liu et al., 2023, Pan et al., 2024, Levy et al., 2024), and Scientific Research (Mühlbacher et al., 2024, Zhang et al., 2025).
-In game industry, there is growing interest in using LLM agents to enhance user game experiences e.g., more intelligent non-player characters (NPCs), monsters, or companions (NVIDIA, 2025).
-In response, many benchmarks have been proposed to assess the ability of LLMs to play games (Hu et al., 2024b).
+Large Language Model (LLM) agents are revolutionizing various industries (Wang et al., 2024a), and well-established benchmarks play a key role in unlocking their abilities on complex tasks, e.g., Coding (Hendrycks et al., 2021, Zhuo et al., 2024), Web Search (Liu et al., 2023, Pan et al., 2024), and Scientific Research (Mühlbacher et al., 2024, Zhang et al., 2025).
+In the game industry, there is growing interest in using LLM agents to enhance user game experiences, e.g., more intelligent non-player characters (NPCs), monsters, or companions (NVIDIA, 2025).
+Also, games that simulate dynamic and uncertain real-world scenarios can serve as effective test-beds for evaluating agents’ high-level decision-making, i.e., system 2 reasoning (Li et al., 2025).
+In response, many benchmarks have been proposed to assess the ability of LLMs by playing games (Hu et al., 2024b, Ahn et al., 2025).
 While these benchmarks have effectively utilized games to evaluate LLMs’ general capabilities, they exhibit three major limitations: 1) they often rely on text-only games or 2D-grid simulators rather than complex real video games, 2) they offer insufficient assessment of agentic modules, such as self-reflection, memory, and tool use, which are essential to complex gameplay, and 3) they lack fine-tuning datasets necessary to adapt pre-trained LLMs into effective gameplay agents, which significantly hinder the adoption of LLM agents in real-world video games.
@@
 These trajectories encapsulate meta-knowledge on how and when to use the agentic strategies to play various genres of games, leading to more resource-efficient and effective gaming agents.
-Our benchmark offers comprehensive evaluation dimensions, including game score leaderboards, competitive LLM battle arenas, and in-depth analyses of visual input state, agentic strategies, and fine-tuning effects.
-Extensive experiments on Orak with 15 LLMs reveal that (1) proprietary LLMs achieve superior performance across games with significant gaps from open-source LLMs, (2) their performance gap becomes narrow in battle scenarios, (3) proprietary LLMs benefit from extended agentic workflows, while open-source LLMs show limited gains, (4) visual states often hinder gameplay performance yet, and (5) fine-tuning enables effective transfer of gameplay meta-knowledge from larger LLMs to smaller ones, leading to generalization in intra-game, out-of-distribution (OOD) game, and non-game unseen scenarios like math and web interaction.
+Our benchmark offers comprehensive evaluation dimensions, including game score leaderboards, competitive LLM battle arenas, and ablation studies of visual input state, agentic strategies, and fine-tuning effects.
+Extensive experiments on Orak with 15 LLMs reveal that (1) proprietary LLMs achieve superior performance across games with significant gaps from open-source LLMs, (2) their performance gap becomes narrow in battle scenarios, (3) proprietary LLMs benefit from extended agentic workflows, while open-source LLMs show limited gains, (4) models fail to extract sufficient value from visual inputs yet, and (5) fine-tuning enables effective transfer of gameplay meta-knowledge from larger LLMs to smaller ones, leading to generalization in intra-game, out-of-distribution (OOD) game, and non-game unseen scenarios like math and web interaction.
 We believe that Orak not only establishes a foundation for building gaming agents but also serves as a critical benchmark for evaluating general LLMs on realistic, long-horizon decision-making tasks.
@@
 GAMA-Bench (Huang et al., 2024), GameBench (Costarelli et al., 2024), GameArena (Hu et al., 2024a), and SmartPlay (Wu et al., 2023) focus on text-based games, assessing LLMs’ ability to navigate and reason on textual environments.
-Barlog (Paglieri et al., 2024) and LVLM-Playground (Wang et al., 2025) are mainly based on 2D-grid games, such as TicTacToe and Chess, to evaluate the spatial and visual reasoning capabilities of LLMs/VLMs.
+Barlog (Paglieri et al., 2024) and LVLM-Playground (Wang et al., 2025a) are mainly based on 2D-grid games, such as TicTacToe and Chess, to evaluate the spatial and visual reasoning capabilities of LLMs/VLMs.
 Using video games, Cradle (Tan et al., 2024) evaluates VLMs in 1 adventure and 3 simulation games, V-MAGE (Zheng et al., 2025) assesses VLMs on 5 action games, DSGBench (Tang et al., 2025) validates LLMs on 6 strategic games, and LMGame-Bench (Hu et al., 2025) mainly focus on puzzle games.
@@
 | Balrog (Paglieri et al., 2024) | Text/2D-grid | ✗ | 6 | LLM/VLM | ✗ | ✗ |
-| LVLM-Playground (Wang et al., 2025) | 2D-grid | ✗ | 6 | VLM | ✗ | ✗ |
+| LVLM-Playground (Wang et al., 2025a) | 2D-grid | ✗ | 6 | VLM | ✗ | ✗ |
 | Cradle (Tan et al., 2024) | Video | ✗ | 4 | VLM | ✗ | ✗ |
@@
 4 Fine-tuning: Aligning Pre-trained LLMs into Game Agents
-| Data example from Reflection module | Data example from Action module | ||||||||||||
+| Data example from Reflection module | Data example from Action module | ||
+| Analyze Mario’s past action using state difference and provide critiques for improving his action.
+You should only respond in the format as below: Self-reflection (Describe self-reflection here) | | (Retrieve the self-reflection from memory) Analyze the current game state, and decide the best action.
+You should only respond in the format as below: Action Jump Level: n (where n is an integer from 0 to 6) | |
+| | Past Game State Mario at (100,100), Bricks at (120, 100), (120, 150) Current Game State Mario at (100,100), Bricks at (120, 100), (120, 150) | | Game State Mario at (100,100) Position of all objects: - Bricks at (120, 100), (120, 150) |
+| | Self-reflection Mario is blocked by bricks. Jump higher to get past. | | Action Jump Level: 6. |
 We collect the fine-tuning dataset from expert LLMs, e.g., GPT-4o and o3-mini, playing all 12 games in Orak using several agentic modules.
@@
 These results suggest that the optimal agentic strategy may depend on the inherent capability of the LLM.
+See Appendices C-N for more detailed ablation analysis for each game.
 5.5 Effect of Visual Input
@@
 OOD-game generalization evaluates whether fine-tuning on a specific set of games enables the model to act better on unseen games.
-Non-game generalization evaluates whether our fine-tuning set with agent gameplay trajectory helps the model to perform better on non-game tasks like math or web navigation.
+Non-game generalization evaluates whether our gameplay fine-tuning set helps the model to perform better on non-game tasks like math or web navigation.
 For Intra-game and Non-game generalizations, we fine-tune LLMs on all 12 games, while for OOD-game generalization, we split 12 games in to 10 training games and 2 test games, i.e., Super Mario and 2048.
+Additionally, to study the effect of data quality and scaling, we perform fine-tuning on high-score (original), low-score (samples with the lowest game score), and their mixed datasets (doubling the data volume).
 See Appendix P for detailed fine-tuning configurations and unseen scenario details in intra-game generalization.
@@
 This is evident in Baba Is You, where fine-tuned models struggle to construct the winning condition required to score above 20.0.
+Also, since StarCraft II requires complex strategic reasoning to defeat enemies, the strategic patterns learned from the fine-tuning set were not sufficient for the model to win on new maps.
 OOD-game Generalization.
@@
 In contrast, the smaller model, Llama-3.2-1B, shows no observable generalization to non-game tasks, which suggests that the extent of generalization may depend on model capacity.
+Effect of Data Quality and Scaling.
+Table 8(a) shows the effect of data quality and scaling in fine-tuning generalization.
+Overall, high-score dataset alone is the most effective, while the low-score dataset contributes little or even degrades performance in certain cases (e.g., Darkest Dungeon).
+Simply scaling up the dataset by mixing high- and low-score samples does not yield a clear improvement, indicating that data quality, rather than sheer quantity, plays a more critical role in achieving better generalization.
+| Data | SF3 | DarkestD | StarCraft2 | SlaySpire | BabaIsYou |
+| No fine-tuning | 12.011.0 | 87.29.5 | 0.00.0 | 0.00.0 | 20.00.0 |
+| high-score (110K) | 40.07.1 | 92.00.2 | 0.00.0 | 10.71.2 | 20.00.0 |
+| low-score (110K) | 18.016.4 | 40.00.0 | 0.00.0 | 8.72.3 | 20.00.0 |
+| mixed (220K) | 18.014.8 | 40.00.0 | 0.00.0 | 10.71.2 | 20.00.0 |
+| Model | Hard (Original) | Easy | ||
+| Paused | Real-time | Paused | Real-time | |
+| gpt-4o-mini (19.6 sec) | 75.00.0 | 0.00.0 | 100.00.0 | 0.00.0 |
+| gpt-4o (27.2 sec) | 100.00.0 | 0.00.0 | 100.00.0 | 100.00.0 |
+| gemini-2.5-pro (99.2 sec) | 100.00.0 | 0.00.0 | 100.00.0 | 0.00.0 |
+5.7 Latency-Aware Evaluation in Real-Time Gameplay
+Among the 12 games in Orak, 3 games, including Street Fighter III, Super Mario, and Starcraft II, require pausing for evaluation.
+Here, we conduct a latency-aware evaluation on StarCraft II by enabling real-time gameplay.
+As shown in Table 8(b), when switching from the paused setting to the real-time mode, all models failed (0 score) under the hard level (the original setup), indicating that timing-sensitive games are strongly affected by inference latency.
+When the difficulty level was reduced to easy, all models achieved perfect scores under the paused setting, but only GPT-4o maintained perfect performance under real-time conditions.
+The average response latencies per step with the reflection–planning agent were 19.6 s for GPT-4o-mini, 27.2 s for GPT-4o, and 99.2 s for Gemini-2.5-pro.
+These results suggest that GPT-4o provides the best trade-off between latency and gameplay accuracy among the three evaluated models.
 6 Discussion
@@
 We leave a complementary direction, providing the full in-game states with rich uncurated texts to LLMs, for future work.
-Further discussions, including cost and licensing, are provided in Appendix R.
+Further discussions, including cost and licensing, are provided in Appendix U.
 Appendix A Game Genre Categorization
@@
 However, a critical aspect in fighting games is understanding the relative positions of the characters, which is not directly provided by the Diambra environment.
-To address this limitation, we employ a lightweight YOLOv11 object detection model (Khanam & Hussain, 2024) to extract the relative positions of the two characters from the game frame.
+To address this limitation, we employ a lightweight YOLOv11 object detection model (Khanam and Hussain, 2024) to extract the relative positions of the two characters from the game frame.
 Based on the computed distance, we classify the spatial relationship into three discrete categories—very close, close, and far—and incorporate this information into the agent’s user prompt.
@@
 The resulting win rate matrix and Elo scores are presented in Figure 4(a).
-We adopt a Bradley-Terry model formulation (Bradley & Terry, 1952) for Elo estimation, where each model’s rating is iteratively optimized using gradient ascent on the log-likelihood of observed outcomes.
+We adopt a Bradley-Terry model formulation (Bradley and Terry, 1952) for Elo estimation, where each model’s rating is iteratively optimized using gradient ascent on the log-likelihood of observed outcomes.
 The gradient is computed based on the expected win probabilities derived from current ratings, using the standard Elo transformation:
@@
 C.5 Result for Street Fighter III
-All reported results in Tables 11–12 are computed as the mean and standard deviation over three to five independent runs, using a ‘zero-shot’ agent for each model configuration.
+All reported results in Tables 12–13 are computed as the mean and standard deviation over five independent runs, using a ‘zero-shot’ agent for each model configuration.
 To establish a baseline for comparison, we additionally included a random agent that selects an arbitrary action uniformly at random.
@@
 This scoring method captures not only victory but also the margin of the win.
-Using these difference-based scores, we then compute Elo ratings for all agents following the Bradley-Terry model (Bradley & Terry, 1952), following the approach used in StreetFighter III multi-agent evaluation in Section C.3.
+Using these difference-based scores, we then compute Elo ratings for all agents following the Bradley-Terry model (Bradley and Terry, 1952), following the approach used in StreetFighter III multi-agent evaluation in Section C.3.
 The resulting army supply difference matrix and Elo scores are presented in Figure 4(b).
@@
 N.5 Result for 2048
-All reported results in Tables 45–45 are computed as the mean and standard deviation over five independent runs, using a ‘zero-shot’ agent for each model configuration.
+All reported results in Tables 46–46 are computed as the mean and standard deviation over five independent runs, using a ‘zero-shot’ agent for each model configuration.
 To establish a baseline for comparison, we additionally included a random agent that selects one of the four possible actions (up, down, left, right) uniformly at random.
@@
 This implies our fine‑tuning set, constructed from a diverse set of games across genres, contains knowledge broadly beneficial for gameplay/decision-making and supports cross‑genre generalizability.
-Appendix R Limitations and Future Work
+Appendix R Unified Performance Comparison
+| Model | Action | Adventure | RPG | Simulation | Strategy | Puzzle | ||||||
+| SF3 | SuperMario | AceAttorney | HerStory | Pokémon | DarkestD | Minecraft | Stardew | StarCraft2 | SlaySpire | BabaIsYou | 2048 | |
+| Main Benchmark Results (Text-only, Default Agent) | ||||||||||||
+| Llama-3.2-1B | 0.00.0 | 18.78.6 | 1.32.2 | 2.11.2 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 6.711.5 | 0.00.1 |
+| Llama-3.2-3B | 13.35.8 | 31.810.1 | 4.61.3 | 4.21.1 | 0.00.0 | 47.539.2 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.30.2 |
+| Qwen-2.5-3B | 20.00.0 | 23.414.1 | 20.017.4 | 1.21.1 | 0.00.0 | 44.822.2 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 13.311.5 | 0.10.1 |
+| Qwen-2.5-7B | 16.711.5 | 27.29.6 | 9.30.2 | 8.52.0 | 0.00.0 | 88.82.0 | 0.00.0 | 0.00.0 | 0.00.0 | 5.00.0 | 20.00.0 | 0.60.4 |
+| Minitron-4B | 16.711.5 | 24.46.0 | 35.74.5 | 4.62.3 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.10.0 |
+| Minitron-8B | 23.35.8 | 31.312.8 | 29.93.6 | 8.21.8 | 0.00.0 | 63.830.4 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.70.7 |
+| Llama-3.3-70B | 33.328.9 | 32.011.8 | 53.923.4 | 46.410.8 | 16.714.4 | 87.28.2 | 0.00.0 | 40.937.2 | 0.00.0 | 5.00.0 | 20.00.0 | 1.41.3 |
+| Qwen2.5-72B | 26.75.8 | 29.98.7 | 14.98.5 | 41.01.7 | 27.89.6 | 84.06.6 | 0.00.0 | 18.016.7 | 0.00.0 | 9.05.3 | 20.00.0 | 0.20.1 |
+| GPT-4o-mini | 16.711.5 | 28.88.8 | 28.42.8 | 21.25.6 | 0.00.0 | 81.35.8 | 46.07.0 | 16.127.8 | 75.050.0 | 3.32.9 | 13.311.5 | 1.11.0 |
+| GPT-4o | 29.714.3 | 34.114.2 | 85.31.5 | 64.25.2 | 38.99.6 | 93.41.5 | 71.07.0 | 81.44.8 | 100.00.0 | 23.622.1 | 20.00.0 | 5.61.5 |
+| GPT-5 | 19.511.0 | 31.610.5 | 59.126.5 | 74.91.7 | 88.94.8 | 92.90.7 | 73.07.0 | 92.38.6 | 25.050.0 | 26.219.4 | 100.00.0 | 10.22.1 |
+| o3-mini | 33.315.3 | 34.914.6 | 91.71.5 | 66.53.6 | 0.00.0 | 89.02.1 | 75.00.0 | 55.116.0 | 25.050.0 | 15.00.0 | 73.346.2 | 25.37.3 |
+| Gemini-2.5-pro | 13.311.5 | 38.014.4 | 55.73.4 | 67.53.3 | 83.30.0 | 93.71.6 | 75.00.0 | 59.210.1 | 100.00.0 | 51.931.9 | 73.346.2 | 5.12.5 |
+| Claude-3.7 | 16.711.5 | 31.78.2 | 81.91.6 | 62.92.6 | 63.919.2 | 89.92.5 | 75.00.0 | 53.620.9 | 50.057.7 | 15.00.0 | 46.746.2 | 5.32.7 |
+| Deepseek-R1 | 20.00.0 | 28.713.2 | 83.31.5 | 67.23.9 | 75.00.0 | 91.71.1 | 41.70.0 | 66.111.5 | 50.057.7 | 24.917.1 | 20.00.0 | 11.53.4 |
+| Agentic Module Ablation (LLaMA-3B / GPT-4o) | ||||||||||||
+| LLaMA-3B (Zeroshot) | 13.35.8 | 21.28.2 | 5.73.2 | 4.21.1 | 0.00.0 | 47.539.2 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.30.2 |
+| LLaMA-3B (Reflection) | 30.017.3 | 32.48.6 | 4.61.3 | 4.41.3 | 0.00.0 | 47.339.0 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.00.0 |
+| LLaMA-3B (Planning) | 20.00.0 | 27.08.4 | 4.61.3 | 5.21.0 | 0.00.0 | 56.323.6 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.10.1 |
+| LLaMA-3B (Ref-Plan) | 16.720.8 | 31.810.1 | 3.80.0 | 5.40.4 | 0.00.0 | 57.031.6 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 20.00.0 | 0.10.2 |
+| GPT-4o (Zeroshot) | 29.714.3 | 29.69.2 | 49.91.3 | 64.25.2 | 33.30.0 | 93.41.5 | 0.00.0 | 34.529.9 | 50.057.7 | 24.79.2 | 20.00.0 | 5.61.5 |
+| GPT-4o (Reflection) | 23.320.8 | 32.315.5 | 85.31.5 | 61.50.4 | 36.14.8 | 85.210.3 | 50.00.0 | 15.64.5 | 0.00.0 | 41.319.6 | 20.00.0 | 3.52.9 |
+| GPT-4o (Planning) | 30.026.5 | 29.412.5 | 52.70.5 | 59.45.7 | 33.30.0 | 82.08.6 | 13.00.0 | 54.920.2 | 50.057.7 | 35.39.2 | 20.00.0 | 6.05.5 |
+| GPT-4o (Ref-Plan) | 23.320.8 | 34.114.2 | 52.80.5 | 62.04.9 | 38.99.6 | 91.62.5 | 50.00.0 | 81.44.8 | 100.00.0 | 36.025.5 | 20.00.0 | 7.05.7 |
+| Modality Ablation (Text / Image / Both) | ||||||||||||
+| Qwen2.5-VL-7B (Text) | 0.00.0 | 26.08.2 | 10.00.0 | 3.21.5 | 2.84.8 | 82.71.5 | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 13.311.5 | 0.10.1 |
+| Qwen2.5-VL-7B (Image) | 3.35.8 | 25.110.1 | – | – | – | – | – | 0.00.0 | – | – | 0.00.0 | 0.20.3 |
+| Qwen2.5-VL-7B (Both) | 3.35.8 | 25.68.0 | 17.613.1 | 2.50.8 | 2.84.8 | 81.83.0 | – | 0.00.0 | 0.00.0 | 0.00.0 | 0.00.0 | 0.40.5 |
+| Qwen2.5-VL-32B (Text) | 16.75.8 | 32.010.2 | 71.921.5 | 15.42.8 | 27.89.6 | 89.91.5 | 0.00.0 | 15.222.7 | 0.00.0 | 0.00.0 | 20.00.0 | 0.20.3 |
+| Qwen2.5-VL-32B (Image) | 33.320.8 | 25.810.7 | – | – | – | – | – | 0.00.0 | – | – | 13.311.5 | 0.40.3 |
+| Qwen2.5-VL-32B (Both) | 33.315.3 | 29.76.6 | 68.610.0 | 17.04.6 | 25.014.4 | 91.03.0 | – | 26.913.9 | 0.00.0 | 0.00.0 | 20.00.0 | 0.20.3 |
+| GPT-4o (Text) | 29.714.3 | 34.114.1 | 85.31.5 | 64.25.2 | 38.99.6 | 93.41.5 | 710.0 | 81.44.8 | 100.00.0 | 23.622.1 | 20.00.0 | 5.61.5 |
+| GPT-4o (Image) | 23.715.9 | 27.113.7 | – | – | – | – | – | 0.00.0 | – | – | 6.711.5 | 1.81.1 |
+| GPT-4o (Both) | 24.314.5 | 27.110.2 | 53.51.7 | 40.629.5 | 41.78.3 | 92.23.0 | – | 41.922.1 | 50.057.7 | 23.622.1 | 20.00.0 | 5.44.5 |
+| Gemini-2.5 (Text) | 13.311.5 | 38.013.4 | 55.73.4 | 67.53.3 | 83.30.0 | 93.71.6 | 75.00.0 | 59.210.1 | 100.00.0 | 51.931.9 | 73.346.2 | 5.12.5 |
+| Gemini-2.5 (Image) | 16.711.5 | 28.510.7 | – | – | – | – | – | 7.69.0 | – | – | 20.00.0 | 5.52.4 |
+| Gemini-2.5 (Both) | 20.010.0 | 40.99.6 | 52.60.8 | 64.92.4 | 83.30.0 | 92.21.8 | – | 60.06.0 | 100.00.0 | 26.219.4 | 86.723.1 | 3.12.6 |
+| Claude-3.7 (Text) | 16.711.5 | 28.713.2 | 81.91.6 | 62.92.6 | 63.919.2 | 89.92.5 | 75.00.0 | 53.620.9 | 50.057.7 | 15.00.0 | 46.746.2 | 5.32.7 |
+| Claude-3.7 (Image) | 23.311.5 | 25.66.4 | – | – | – | – | – | 0.00.0 | – | – | 20.00.0 | 8.44.0 |
+| Claude-3.7 (Both) | 33.35.8 | 22.66.3 | 71.317.3 | 63.63.1 | 72.24.7 | 90.15.7 | – | 49.81.0 | 50.057.7 | 9.74.6 | 20.00.0 | 6.70.9 |
+| Human Novice | ||||||||||||
+| Human (Novice) | 20.020.0 | 100.00.0 | 87.82.6 | 72.62.8 | 86.112.7 | 90.32.0 | 70.80.0 | 69.819.9 | 33.357.7 | 52.923.2 | 100.00.0 | 22.710.4 |
+Table 50 compares the overall LLM/VLM agent results across agentic modules and modalities. Interestingly, one of the LLM agent that are using text-only states performs the best for each game.
+Appendix S Statistical Significance Analysis
+| Model | Metric | SF3 | SuperMario | AceAttorney | HerStory | Pokémon | DarkestD | Minecraft | Stardew | StarCraft2 | SlaySpire | BabaIsYou | 2048 |
+| – | # seed | 5 | 20 | 3 | 3 | 3 | 3 | 3 | 3 | 4 | 3 | 3 | 5 |
+| GPT-4o | mean | 29.7 | 34.1 | 85.3 | 64.2 | 38.9 | 93.4 | 71.0 | 81.4 | 100.0 | 23.6 | 20.0 | 5.6 |
+| std | 14.3 | 14.2 | 1.5 | 5.2 | 9.6 | 1.5 | 7.0 | 4.8 | 0.0 | 22.1 | 0.0 | 1.5 | |
+| CI (95%) | 17.76 | 6.65 | 3.73 | 12.92 | 23.85 | 3.73 | 17.39 | 11.92 | 0.00 | 54.90 | 0.00 | 1.86 |
+This section reports three complementary statistical analyses: (1) per-game confidence intervals, (2) per-game significance tests using Welch’s t-test (Welch, 1947), and (3) across-game paired t-tests (Fisher, 1925).
+Per-Game Confidence Intervals. For each game, we report the sample mean , sample standard deviation , the number of seeds , and the two–sided confidence interval half–width (“CI (95%)”) computed as
+Table 51 summarizes the results for GPT-4o.
+| Game | t-stat | p-value | Interpretation |
+| SF3 | 4.644 | 0.0097 | significant |
+| Super Mario | 4.149 | 0.0002 | significant |
+| Ace Attorney | 54.641 | significant | |
+| Her Story | 20.155 | 0.0015 | significant |
+| Pokémon | 7.018 | 0.0197 | significant |
+| Darkest Dungeon | 107.849 | 0.0001 | significant |
+| Minecraft | 17.568 | 0.0032 | significant |
+| Stardew Valley | 29.373 | 0.0012 | significant |
+| StarCraft II | – | – | deterministic |
+| Slay the Spire | 1.850 | 0.2056 | not significant |
+| Baba Is You | 2.003 | 0.1831 | not significant |
+| 2048 | 8.329 | 0.0011 | significant |
+| Game | t-stat | p-value | Interpretation |
+| SF3 | 0.435 | 0.681 | not significant |
+| Super Mario | 1.128 | 0.268 | not significant |
+| Ace Attorney | 14.127 | 0.0039 | significant |
+| Her Story | 7.345 | 0.0104 | significant |
+| Pokémon | 1.416 | 0.230 | not significant |
+| Darkest Dungeon | 2.406 | 0.126 | not significant |
+| Minecraft | 17.568 | 0.0032 | significant |
+| Stardew Valley | 6.320 | 0.0164 | significant |
+| StarCraft II | – | – | deterministic |
+| Slay the Spire | 1.113 | 0.371 | not significant |
+| Baba Is You | – | – | deterministic tie |
+| 2048 | 8.032 | 0.0013 | significant |
+Per-Game Significance Tests.
+We compare GPT-4o against Llama-3.2-3B and Qwen-2.5-72B on each game using Welch’s t-test (Welch, 1947).
+As shown in Table 52, GPT-4o and Llama-3.2-3B shows significant or deterministic differences in 10 out of 12 games, while GPT-4o and Qwen-2.5-72B shows significant or deterministic differences in 7 out of 12 games.
+Here, “Deterministic” indicates both models have zero empirical variance and unequal means; “deterministic tie” indicates both zero variance and equal means.
+| Comparison | t-stat | p-value | Interpretation |
+| GPT-4o vs. Llama-3.2-1B | 5.227 | 0.00028 | significant |
+| GPT-4o vs. Qwen-2.5-72B | 3.122 | 0.00971 | significant |
+Appendix T Discussion for Integrating Orak with Multi-Turn RL Fine-Tuning Frameworks
+Orak can be readily integrated with existing multi-turn RL fine-tuning frameworks for LLM agents (Wang et al., 2025b, Zeng et al., 2025).
+In game-based environments, rewards are often sparse, being provided only at the end of an episode (e.g., final score).
+Therefore, careful reward design and dynamic data extraction are crucial for stable and effective optimization.
+Reward Design. We discuss two types of reward designs.
+(1) Reward Discounting. Following common practices in reinforcement learning, we apply reward discounting to estimate intermediate rewards during gameplay:
+where denotes the discount factor and is the final game score.
+(2) Turn-Based Rewards (when available). For games that provide intermediate feedback, such as 2048 or Super Mario, a turn-level reward can be incorporated as:
+where is a balancing coefficient and is the turn-level score. This formulation allows smoother credit assignment across gameplay turns.
+Dynamic Data Extraction.
+During fine-tuning, the model repeatedly performs data rollouts, reward estimation, and multi-turn RL optimization across multiple episodes.
+As the policy improves, it naturally encounters increasingly challenging game scenarios, promoting progressive data extraction and richer experience replay.
+After each episode, validation accuracy is compared between the current and previous policy models, and the superior one is retained.
+This iterative selection procedure contributes to stable and robust policy improvement over time.
+Appendix U Limitations and Future Work
 Cost Consideration.